### v1

Abstract
Artificial Intelligence (AI) has shown great promise in electrocardiogram (ECG) analysis and cardiovascular disease detection. However, developing a general AI-ECG model has been challenging due to inter-individual variability and the diversity of ECG diagnoses, limiting existing models to specific diagnostic tasks and datasets. Moreover, current AI-ECG models struggle to achieve comparable performance between single-lead and 12-lead ECGs, limiting the application of AI-ECG to portable and wearable ECG devices. To address these limitations, we introduce an ECG Foundation Model (ECGFounder), a general-purpose model that leverages real-world ECG annotations from cardiology experts to broaden the diagnostic capabilities of ECG analysis. ECGFounder is trained on over 10 million ECGs with 150 label categories from the Harvard-Emory ECG Database, enabling comprehensive cardiovascular disease diagnosis through ECG analysis. The model is designed to be both effective out-of-the-box and fine-tunable for downstream tasks, maximizing usability. More importantly, we extend its application to single-lead ECGs, enabling complex condition diagnoses and supporting various downstream tasks in mobile and remote monitoring scenarios. Experimental results demonstrate that ECGFounder achieves expert-level performance on internal validation sets for both 12-lead and single-lead ECGs, while also exhibiting strong classification performance and generalization across various diagnoses on external validation sets. When fine-tuned, ECGFounder outperforms baseline models in demographics detection, clinical event detection, and cross-modality cardiac rhythm diagnosis. ECGFounder not only enhances existing cardiovascular diagnostics capabilities but also facilitates integration into mobile and remote monitoring devices, enabling continuous, real-time cardiac health monitoring crucial for early detection and management of chronic heart conditions. The trained model and data will be publicly released upon publication through the Brain Data Science Platform (bdsp.io), promoting further research and development in this field. Our code is available at https://github.com/bdsp-core/ECGFounder.
keywords:
[…]
1 Introduction
Electrocardiogram (ECG) is a crucial diagnostic tool that measures and records the electrical activity of the heart using electrodes placed on the skin.[2] ECG recordings are vital diagnosis and monitoring of cardiac health conditions.[14] In addition to common arrhythmias, ECGs can be used to diagnose conditions such as myocardial ischemia, myocardial hypertrophy, and myocardial infarction.[18] However, fully interpreting an ECG is complex and difficult. A typical ECG expert undergoes nearly 10 years of training, including medical school, internal medicine residency, and specialized ECG training.[2] Despite this, comprehensive and accurate interpretation of ECGs remains a challenge even for qualified experts. In real clinical scenarios an ECG may present multiple diagnoses simultaneously. For instance, a patient with coronary heart disease might also exhibit bundle branch block, ST-segment depression, and T-wave inversion. Even ECG experts may occasionally overlook or mislabel findings due to human error. There is an urgent need for an AI-powered automated model capable of identifying all ECG abnormalities.
Traditional ECG analysis focuses on extracting waveform and rhythm features from ECG signals for arrhythmia classification and ECG anomaly detection.[9, 28] These methods are simple and effective, but limited, as they rely on expert knowledge and prior experience, which may fail to capture complex patterns in the data. Manual feature selection can lead to missing important information or overfitting. To address this limitation, some studies have introduced signal analysis techniques, such as Fourier transform and wavelet transform, which allow for automatic feature extraction and classification of ECG signals. [32, 6] Compared to traditional ECG analysis, signal analysis methods offer higher accuracy. However, these models face challenges in handling and analyzing large-scale ECG data, and they can be computationally intensive.
Recent work has integrated deep learning models into ECG analysis, significantly improving diagnostic accuracy and efficiency.[42] Current deep learning approaches in ECG interpretation primarily employ Convolutional Neural Networks (CNNs)[16, 43, 25], Recurrent Neural Networks (RNNs)[34], Transformers[24] and their variants, which have proven effective in identifying common arrhythmias and conduction blocks.[10, 1, 30]
However, due to limitations of existing ECG databases in terms of sample size, patient numbers, and the variety of diagnoses, these models have not yet addressed the challenges of inter-individual variability and the diversity of diagnoses[23]. This restricts the models to performing only a limited set of diagnostic tasks and datasets. Additionally, current ECG models do not achieve the same level of performance for single-lead ECGs as they do for 12-lead ECGs.[17, 29] These challenges highlight the practical limitations of training an ECG model from scratch on small-scale datasets, making it difficult to extend the model to real-world ECG analysis. A promising approach is to leverage transfer learning and the strong generalization capabilities of foundation models to enhance the performance of ECG models.
Foundation models have recently made significant strides in medical AI. In retinal disease diagnosis, the RETFound model, through pre-training on a large number of retinal images, has achieved excellent performance across various clinical diagnostic tasks.[41] In computational pathology, the UNI model was trained on a vast amount of whole-slide images, reaching expert-level performance in multiple cancer diagnostic tasks.[4] In these studies, foundation models are defined as large-scale AI models trained on extensive datasets, capable of adapting to a wide range of downstream tasks. Specifically, they meet the following criteria: 1) the pre-training dataset is large in scale, 2) the model has an enormous number of parameters, and 3) it can perform a wide range of downstream tasks. Although recent work has made substantial progress in many medical fields, there is currently no foundation model for ECGs.[36, 3]
To fully leverage the advantages of foundation models, we develop an ECG Foundation Model—a groundbreaking approach aimed at revolutionizing ECG analysis. This foundation model, the first of its kind, is capable of diagnosing 150 types of cardiac abnormalities covered by existing ECG classifications by utilizing over ten million clinically annotated real-world ECG datasets. To address the inherent challenges of incomplete annotations in real-world data, we introduce a novel method for pre-processing and training on these annotations, ensuring robust performance even under sub-optimal conditions. Moreover, by training the single-lead ECG model based on lead augmentation, we are able to maintain high diagnostic performance on single-lead ECGs as well. We validate the model’s performance on both internal and external test sets, where it consistently matches expert-level diagnoses. In downstream task fine-tuning, we demonstrate ECGFounder’s versatility in addressing various tasks, including demographics detection, clinical event detection, and cross-modality diagnosis (Figure 1b). Specifically, we evaluate ECGFounder on 12 clinical tasks, such as ECG age regression and classification, sex detection, chronic kidney disease (CKD) detection, chronic heart disease (CHD) detection, regression and abnormal classification of left ventricular ejection fraction (LVEF), regression and abnormal classification of NT-proBNP, and atrial fibrillation detection based on photoplethysmography(PPG). The results of these downstream tasks highlight ECGFounder’s potential as a foundational model for the further development of AI-ECG models.
Figure 1: ECGFounder is a general ECG encoder based on the RegNet architecture. (a) ECGFounder was trained on 10,771,552 ECGs with 150 types of ECG diagnostic labels. Due to the presence of missing diagnoses in real-world expert annotations, we implemented positive label augmentation by modifying the loss function of the pre-training method. For more details, see the Methods section. (b) ECGFounder was applied to 12 clinical downstream tasks, covering ECG diagnosis, demographics inference, clinical event detection, and cross-modality diagnosis. In comparison with baseline methods, ECGFounder achieved state-of-the-art performance across all tasks. (c) A comparison of the HEEDB dataset used by ECGFounder with other large ECG datasets. For more details, see Table LABEL:tb:data. (d) Frequency distribution of HEEDB labels. The x-axis is the number of samples in logarithm, and the y-axis is the label type. The label types are categorized based on their frequency into common, uncommon, and rare sets. The common set has a frequency greater than $10^{5}$. The uncommon set has a frequency greater than $10^{4}$ but less than $10^{5}$. The rare set has a frequency greater than $10^{3}$ but less than $10^{4}$. PPG., photoplethysmography; LVEF., left ventricular ejection fraction; CKD., chronic kidney disease; CHD., chronic heart disease.
2 Methods
2.1 Datasets and pre-processing
Our dataset, the Harvard-Emory ECG Database (HEEDB), is currently the largest open-access ECG dataset, containing 10,771,552 expert-annotated ECGs from 1,818,247 unique subjects.[19] These ECGs are predominantly 10-second, 12-lead clinical ECGs. The dataset includes annotations from cardiologists and ECG technicians paired with the ECGs. These annotations consist of discrete text reports that primarily describe the morphology, rhythm, and diagnostic information of the ECGs. Experts used the Marquette 12SL ECG Analysis Program (GE Healthcare) version 4 to assist with annotations,[12] which provides waveform parameters for doctors’ reference. The program offers many diagnostic categories, allowing doctors to simply click on corresponding category labels on the computer, avoiding manual entry of diagnostic statements. To extract descriptive information about ECGs from these discrete labels, we utilized regular expressions to parse the annotations, tallying each independent label. In total, there were 287 independent phrases. After reviewing with doctors, we removed phrases that were irrelevant to ECG descriptions while retaining meaningful phrases. Ultimately, we defined 150 meaningful labels, including rich diagnostic information as well as specific rhythm and morphological descriptions (See Supplement for more details about labels).
Our external validation data comprised three large ECG databases (DB): The CODE-test DB [31], the PTB-XL DB[38], and PhysioNet Challenge-2017 DB [5]. The CODE-test DB is composed of ECG records from 827 patients across 811 municipalities in Brazil, collected by the Telehealth Network of Minas Gerais (TNMG). There are six common arrhythmia labels in this ECG DB, annotated by several experienced ECG experts.
The PTB-XL dataset contains 21,837 clinical ECGs from 18,885 patients in Germany. Each ECG is a 10-second, 12-lead recording. The labels were reviewed and verified by two physicians. This dataset is currently one of the best publicly accessible ECG collections, both in terms of the number of samples and the quality of the labels.[38]
The PhysioNet Challenge-2017 is a large single-lead ECG dataset. The ECG recordings were collected using the AliveCor device.[5] The training set includes 8,528 single-lead ECG recordings, with durations ranging from 9 to 60 seconds, while the test set contains 3,658 ECG recordings of similar lengths. The dataset requires classification of ECG recordings into normal rhythm, atrial fibrillation rhythm, other rhythms, and noise.
Additionally, we utilized the MIMIC-IV-ECG DB to fine-tune our model for various downstream tasks. The MIMIC-IV-ECG dataset is part of the MIMIC series, focusing on the collection and analysis of ECG data.[7] MIMIC-IV-ECG originates from real clinical settings at Beth Israel Deaconess Medical Center (BIDMC) in Boston, USA, and contains a large number of clinical ECG recordings from patients treated in the intensive care unit (ICU). Moreover, ECG recordings in the dataset can be matched with the electronic health records of patients in the MIMIC-ED, allowing ECG data to be associated with specific conditions. Here, we linked several clinical downstream tasks, such as age, sex, chronic kidney disease (CKD), chronic heart disease (CHD), left ventricular ejection fraction (LVEF) and NT-proBNP with ECG data to explore the performance improvement of the fine-tuned ECGFounder model in detecting other diseases and clinical events. More details about the split and labels of MIMIC-IV-ECG DB can be found in Supplement.
To explore ECGFounder’s generalization capabilities on other similar physiological signals, we fine-tuned and evaluated it using the DeepBeat dataset, a PPG-based atrial fibrillation (AF) detection dataset.[37] The dataset includes over 500,000 labeled signals from more than 100 individuals.
For ECG preprocessing, we applied linear interpolation to resample ECG frequencies to 500Hz. We used a high-pass filter with a cutoff frequency of 1 Hz to suppress residual baseline drift and a second-order 30 Hz Butterworth low-pass filter to reduce high-frequency noise. A 50/60 Hz notch filter was utilized to eliminate electrical interference. From ECG records longer than 10 seconds, we extracted 10-second windows in sequence. If a sequence was less than 10 seconds, we applied zero padding. We normalized all signals using the mean and standard deviation of each individual signal segment before inputting them into the model. If a signal was missing, we assigned it a value of 0.
2.2 Model architecture
To establish an ECG foundation model, we require an architecture that is tailored for ECG, capable of learning generalizable representations from large-scale ECG datasets. The growth in the size of ECG data and the increase in the number of leads mean that the model must not only consider temporal information of the data but also spatial relationships, i.e., interactions between different leads and the overall pattern of cardiac electrical activity. This is crucial for the ECG foundation model to be applicable in real clinical ECG scenarios, to mitigate the impact of non-uniform ECG durations and missing leads in the training dataset.
Considering these factors, we built our model architecture on top of the RegNet architecture.[15, 27] This structure begins with a stage-wise network design where each stage consists of a set number of blocks and channels that scale with network depth. This is beneficial for us to expand and design the blocks and channels of the ECG foundation model. Unlike traditional uniform scaling across layers, RegNet employs a quantized linear model to predict optimal widths and depths, ensuring efficient performance across a range of model sizes. The initial layers focus on capturing low-level features with fewer channels, which gradually increase as the network deepens, thus optimizing computational efficiency and model capacity. Following this, the model utilizes a series of bottleneck blocks that combine group convolutions and channel-wise attention mechanisms, enhancing the richness of representation in both temporal and spatial dimensions while controlling model complexity. This tailored configuration makes the model suitable for ECG data. More details about the model architecture can be found in the Supplement.
2.3 Training with real-world annotations: noisy, imbalanced, positive unlabeled
Unlike conventional ECG diagnostic models that typically use single-label classification methods, we employed multi-label classification during the training phase of the ECG foundation model. This approach aligns more closely with clinical practice, where an ECG diagnosis often consists of multiple diagnostic labels. For example, an abnormal ECG diagnosis might be something like normal sinus rhythm | premature ventricular complexes | premature ectopic complexes. Additionally, multi-label diagnostics better meet the application needs of clinicians. During training, utilizing a multi-label classification approach provides our model with rich semantic value and facilitates generalization to different annotation systems. However, the nature of multi-label annotations means that it is almost impossible to achieve perfect multi-label data, especially in cases like ECG where there are many diagnostic categories. Generally speaking, if a cardiology expert can annotate three classes simultaneously, it is considered excellent. However, the actual number of ECG diagnostic labels far exceeds three classes.[18] To address the challenge brought by incomplete ECG annotations, we introduce positive unlabeled (PU) learning method.
PU learning is defined as when positive samples are present in the data but unlabeled, but the labeled positive samples are correct, meaning that the unlabeled samples are not necessarily negative examples.[40] Typically, PU learning leads to a severe imbalance in the predicted probability distribution: an ECG usually contains a few positive diagnostic labels, with the remaining labels being treated as negative, leading to a severe positive-negative imbalance where the negative samples far outnumber the positive samples for each label. The inherent scarcity of positive examples, combined with the missing labels, exacerbates the imbalance, making it more challenging to learn from the positive examples. When using conventional loss functions, such as the Binary Cross-Entropy (BCE) loss function, the model tends to learn from simpler samples, namely the true negative (TN) samples, while the more challenging samples, namely the false negative (FN) samples, are harder to fit. When the model is biased towards learning TN samples, its predicted probabilities tend to skew towards 0 rather than 1, which diminishes its ability to predict positive samples and weakens its potential to detect clinically important abnormalities.
To enhance the model’s ability to fit representations of positive samples, we improved the multi-label classification loss function, enabling the model to correct missing labels by adjusting the weights of positive and negative samples.[39] Typically, for missing labels, a well-trained multi-label model’s predicted probability should be close to 1 rather than 0. Therefore, by applying a smaller weight to the loss of negative labels with predicted probabilities close to 1, we can mitigate the impact of missing labels. The loss function is given by:
$\mathcal{L}=-(\gamma-p)p^{2}.$ (1)
Here, $\gamma$ is a hyperparameter of the model and $p$ is the predicted probability of model. In this case, it is set to $\gamma=1.5$, which we find optimally balances the weights, allowing the model to learn good representations of both positive and negative samples.
Our model training used AdamW to minimize the loss function, with an initial learning rate set to 0.001.[21] The learning rate decayed by a factor of 10 every 5 epochs. The trainable temperature parameter was initialized to 0. Training spanned a maximum of 20 epochs, with early stopping based on validation loss. We used a batch size of 1024.
2.4 Training a single-lead ECG model based on leads augmentation
The development of portable and wearable ECG devices in recent years has revolutionized continuous cardiac monitoring, offering a non-invasive method for real-time health assessment. Beyond the conventional diagnosis of cardiac arrhythmias, another critical challenge in this field is accurately detecting and interpreting ECG axis deviation on single-lead ECGs from portable devices (typically lead I), which can significantly impact the diagnosis of various cardiac abnormalities. For instance, left axis deviation can provide additional insights into diagnoses like left ventricular hypertrophy, left bundle branch block, left anterior fascicular block, pre-excitation syndromes, and inferior myocardial infarction.[17, 8] To address this issue, we developed a novel training method utilizing lead-augmented portable ECG models. By systematically enhancing standard 12-lead ECG data, we simulate various clinical scenarios of axis inversion, thereby enhancing the model’s robustness and versatility. Understanding the relationship between ECG vectors and leads is crucial for this method. The cardiac electrical activity generates a vector representation of cardiac signals, captured by different ECG leads placed on the body. Each lead provides a unique perspective of the cardiac electrical axis, offering a comprehensive view when combined. The standard 12-lead ECG system includes limb leads (I, II, III, aVR, aVL, aVF) and precordial leads (V1-V6), with each lead representing a specific projection of the cardiac electrical vector.[22] Specifically, we primarily utilize the ECG signals from limb leads. Based on the angular position of each limb lead’s axis relative to the heart, we consider lead I as the center of the semicircle (i.e., 0°) and calculate the signals from all leads around the semicircle (i.e., from -90° to 90°), thereby obtaining six additional leads (aVL, -aVR, II, -III, aVF, -aVF). We then trained a model for portable ECG devices, extracting the lead I ECG from the HEEDB 12-lead data and randomly incorporating one of the remaining six augmented leads into the model for training with a 50% probability. This ensures that the model can learn arrhythmia features from lead I data and axis deviation from the additional six leads. Additionally, we have scaled down the model’s parameter size to optimize for portable devices with limited computational resources.
2.5 Fine-tuning on ECG foundation model
When adapting to specific ECG downstream tasks, we need to retain the parameters of the base model and discard the initial classification linear layer. The number of classes in the downstream task determines the number of neurons needed in the final layer of the new linear layer. The training objective is to generate classification outputs that match the labels. To fully leverage the potential of the foundation model, we experimented with two fine-tuning strategies. The first fine-tuning strategy involved loading the pre-trained model parameters and fine-tuning the entire model. The second fine-tuning strategy involved loading the model parameters, freezing the encoder’s parameters (i.e., not altering the encoder parameters), and only fine-tuning the final linear layer of the model.
The batch size is 256. The total training period is 30 epochs, with a learning rate adjustment strategy that utilizes the scheduler. After every epoch, the scheduler monitors the specified metric, and if the performance does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 0.1. The learning rate reduction is triggered based on the maximization of the monitored metric. This approach ensures a dynamic adjustment of the learning rate depending on the training progress. After training in each epoch, the model is evaluated on the validation set.[4] The model weights with the highest AUC on the validation set are saved as a checkpoint for future evaluation.
2.6 Clinical validation
To validate and compare the performance of our model, we established a committee consisting of three experienced ECG physicians to annotate the internal test set. We developed an ECG annotation system for physicians to use, and set up 20 types of labels including cardiac rate abnormalities, conduction blocks, myocardial dilation, myocardial infarction, and ECG morphologies, with several sub-labels under each category. The Supplement displays the complete list of label types. After independent annotation by the committee, a consensus determination was made; ECGs that did not reach consensus were removed, providing an expert standard for model evaluation. ECGs that the committee could not interpret or agree upon were eliminated from our test dataset.
Additionally, to compare the diagnostic accuracy between the model and experts, five additional ECG physicians were involved in this study. They were given specific instructions on how to use the system. The Supplement shows the doctors’ ages, levels of experience, and education. Each doctor was required to independently annotate each ECG from the internal test set. The annotations from these physicians were then compared to the model’s results.
2.7 Statistical analysis
The evaluation of the model was conducted by calculating accuracy, the area under the Receiver Operating Characteristic (ROC) curve (AUC), sensitivity, specificity, and the F1 score, each with bilateral 95% confidence intervals (CI).
3 Results
Our dataset comprises 10,771,552 expert-annotated ECGs recorded from 1,818,247 unique subjects using 12-lead ECG machines, covering the period from the 1990s to present, with a recording duration of 10 seconds. The average age of patients is 60.6 years, with 48.7% being female.[19] After annotation cleaning, the training dataset became a multi-label dataset, with over 10,558,162 (98%) of ECGs being multi-labeled. The development dataset includes 7,519,035 ECGs from 1,725,643 patients, and the test dataset includes 834,926 ECGs from 265,818 patients. The committee internal test set includes 523 ECGs recorded between January 10, 2023, and January 20, 2023, corresponding to 523 unique patients. We obtained a total of 150 labels describing the ECGs, including heart rate abnormalities, conduction abnormalities, cardiac hypertrophy, myocardial infarction, myocardial ischemia, pacemaker devices, and ECG waveform morphology. Using these labels, we trained two deep learning ECG models separately: one for standard 12-lead ECGs and the other designed for single-lead portable device ECGs (lead I). We evaluated the performance of both models on internal and external test sets.
We first conducted large-scale validation of ECGFounder using an internal test dataset comprising 834,926 ECGs, to assess its consistency with diagnoses confirmed by clinically recognized cardiologists (Figure 2). ECGFounder demonstrated an impressive AUC of at least 0.950 across 82 diagnostic categories and at least 0.900 across 112 categories. Overall, the AUCs were particularly high for categories such as rhythm disorders, conduction abnormalities, cardiac hypertrophy, and myocardial infarction. In contrast, the AUCs were lower for categories involving waveform changes, which may be attributed to the annotation habits of certain physicians who tend to place less emphasis on documenting waveform changes in their diagnoses.
We validated the deep learning model for 12-lead ECGs using the committee internal test set, which includes 523 ECGs recorded between January 10, 2023, and January 20, 2023, corresponding to 523 unique patients (49% female; mean age 57.2 years). In internal test set experiments for 12-lead ECGs, the algorithm’s average AUC score for diagnosing all 20 classifications was 0.968 (95% CI, 0.965–0.969), the PRAUC score was 0.912 (95% CI, 0.910–0.914), sensitivity was 0.971 (95% CI, 0.970–0.973), and specificity was 0.937 (95% CI, 0.935–0.938). When comparing the model with cardiologists, our model achieved an overall average F1 score of 0.683, outperforming the cardiologists’ average F1 score of 0.640. The model’s performance was compared with cardiologists’ performance across 20 diagnostic categories (Table LABEL:tab:internal). The model’s F1 scores were higher than those of the cardiologists in 12 categories (atrial fibrillation, atrial flutter, incomplete right bundle branch block, lateral infarct, left bundle branch block, normal sinus rhythm, premature atrial complexes, right axis deviation, right bundle branch block, sinus bradycardia, sinus rhythm, with sinus arrhythmia). Notably, the model’s F1 scores for atrial flutter, incomplete right bundle branch block, lateral infarct, left bundle branch block, and right axis deviation were significantly higher than those of the cardiologists. The model’s F1 scores for anterior infarct, atrial-paced rhythm, inferior infarct, and ventricular tachycardia were slightly lower than those of the cardiologists. When comparing the model’s ROC curve with the true positive rate and false positive rate of cardiologists, the model outperformed the average performance of cardiologists for most labels.
In the external test set experiments for 12-lead ECGs, we evaluated the performance of our model and other models on the CODE-test and PTB-XL datasets. On CODE-test, our model achieved an average AUC of 0.981, outperforming other baseline models such as CinC2020-CTN, which had an average AUC of 0.963, and 0.980 (S12L-ECG)(Table 1). On PTB-XL, because the other two models can only diagnose arrhythmias and cannot complete other diagnostic classifications in the broader classes on PhysioNet Challenge-2020, we validated this dataset only on the class which the models have. Our model achieved an average F1 score of 0.732 and an average AUC of 0.924(Table 1). These results indicate that our model generalizes to different regions, hospitals, and patients.
In the internal test set experiments focused on single-lead ECGs, the model demonstrated excellent performance in rhythm-type diagnoses(Figure 2). Notably, it achieved an AUC above 0.95 for common heart rate abnormalities such as normal sinus rhythm, sinus bradycardia, sinus tachycardia, marked sinus bradycardia, sinus arrhythmia, marked sinus arrhythmia, and atrial fibrillation. These diagnoses can be reliably identified from single-lead ECGs. Additionally, the model achieved an AUC above 0.8 for diagnoses including premature ventricular complexes, premature supraventricular complexes, pacemaker, 1st degree AV block, branch block, fascicular block, and chamber enlargement. It is worth noting that for some of these diagnoses, this is the first time such exceptional performance has been observed using single-lead ECGs. In the diagnosis of myocardial infarctions, the model also showed good diagnostic performance for lateral infarct, anterolateral infarct, acute MI, and STEMI. In ECG research, these diagnoses are associated with shifts in the heart’s electrical axis. This suggests that data augmentation methods based on the electrical axis are effective for training single-lead ECG models.
In the external test set experiments on single-lead ECG devices, our model accurately classified sinus rhythm and atrial fibrillation. The model’s performance, as shown in the table, achieved an AUC of 0.975 and 0.957 for normal sinus rhythm and atrial fibrillation, respectively (Table 10). This indicates that the model achieves excellent performance on portable and wearable ECGs in real-world scenarios.
Figure 2: Results of AUROC on 12-lead and single-lead ECG in internal test set.
Figure 3: (a) Results of NTproBNP classification, NTproBNP regression, CKD classification, LVEF classification, LVEF regression and CHD classification tasks of ECGFounder and other baseline models for clinical events detection tasks. (b) Results of age classification over 65 years old, age regression and sex classification tasks of ECGFounder and other baseline models for demographics detection tasks. (c) Results of PPG atrial fibrillation detection tasks.
We next validated the performance of ECGFounder in transfer learning. Our model was fine-tuned and evaluated on six downstream tasks using supervised learning on the MIMIC-IV-ECG dataset, resulting in both a 12-lead ECG model and a single-lead ECG model. The fine-tuning results are shown in Figures 3. As shown, ECGFounder outperforms the baseline methods in every downstream task. Specifically, it achieves 2 to 3 percentage points higher performance than ECG-SimCLR and 4 to 6 points higher than ECG-ResNet in age, sex, NTproBNP, CKD, and CHD detection. Additionally, to compare with the previously published CKD study [13], we created an independent test set consisting of individual patients from the MIMIC-IV-ECG dataset for validation. As shown in the results (Table 2), our model surpasses the previously studied model in all six classifications for both 12-lead and single-lead ECGs.
Table 1: Performance of other ECG deep learning model and ECGFounder on external test set CODE-test and PTB-XL.[31, 38] In particular, for the baseline method S12L-ECG, the CODE-test is the internal test set.
Dataset CODE-test Models ECGFounder S12L-ECG(Internal)[31] CTN[24] ECG-SE-ResNet[43] Count AUC Sens Spec AUC Sens Spec AUC Sens Spec AUC Sens Spec SINUS BRADYCARDIA 16 0.967 1.000 0.955 0.955 0.938 0.996 0.965 0.987 0.942 0.932 0.995 0.937 ATRIAL FIBRILLATION 13 0.999 1.000 0.996 0.963 0.769 1.000 0.966 0.972 0.969 0.976 0.980 0.970 SINUS TACHYCARDIA 37 0.989 0.974 0.970 0.977 0.973 0.997 0.972 0.958 0.943 0.976 0.950 0.957 RIGHT BUNDLE BRANCH BLOCK 34 0.989 0.971 0.971 0.995 1.000 0.995 0.989 0.946 0.986 0.965 0.949 0.947 LEFT BUNDLE BRANCH BLOCK 30 0.998 1.000 0.996 1.000 1.000 1.000 0.961 0.988 0.942 0.971 0.949 0.976 WITH 1ST DEGREE AV BLOCK 28 0.949 0.864 0.957 0.989 0.929 0.995 0.930 0.862 0.925 0.958 0.856 0.922 Dataset PTB-XL Models ECGFounder S12L-ECG[31] CTN[24] ECG-SE-ResNet[43] ANTERIOR INFARCT 36 0.635 0.816 0.447 / / / / / / / / / ANTEROLATERAL INFARCT 29 0.945 0.837 0.918 / / / / / / / / / ANTEROSEPTAL INFARCT 2415 0.945 0.891 0.882 / / / / / / / / / ATRIAL FIBRILLATION 1514 0.993 0.969 0.982 0.925 0.776 0.989 0.953 0.935 0.915 0.950 0.946 0.907 ATRIAL FLUTTER 73 0.993 0.973 0.969 / / / 0.948 0.962 0.937 0.941 0.966 0.956 ELECTRONIC ATRIAL PACEMAKER 298 0.909 0.935 0.745 / / / / / / / / / INFERIOR INFARCT 276 0.851 0.825 0.732 / / / / / / / / / LATERAL INFARCT 1066 0.916 0.819 0.859 / / / / / / / / / LEFT ANTERIOR FASCICULAR BLOCK 167 0.972 0.922 0.919 / / / 0.965 0.882 0.914 0.968 0.819 0.901 LEFT ATRIAL ENLARGEMENT 427 0.799 0.742 0.704 / / / / / / / / / LEFT BUNDLE BRANCH BLOCK 539 0.989 0.969 0.965 0.983 0.963 0.991 0.912 0.932 0.933 0.931 0.917 0.930 LEFT POSTERIOR FASCICULAR BLOCK 18 0.840 0.605 0.904 / / / / / / / / / LEFT VENTRICULAR HYPERTROPHY 2419 0.900 0.860 0.788 / / / / / / / / / LOW VOLTAGE QRS 182 0.581 0.690 0.423 / / / 0.605 0.605 0.814 0.593 0.655 0.741 N-SPECIFIC INTRAVENTRICULAR BLOCK 78 0.766 0.339 0.932 / / / 0.788 0.487 0.906 0.776 0.699 0.793 NORMAL ECG 987 0.887 0.952 0.639 / / / / / / / / / PREMATURE VENTRICULAR COMPLEXES 1146 0.987 0.965 0.959 / / / 0.963 0.955 0.921 0.937 0.939 0.880 QT HAS LENGTHENED 119 0.931 0.897 0.827 / / / 0.930 0.881 0.957 0.930 0.901 0.930 RIGHT ATRIAL ENLARGEMENT 99 0.959 0.869 0.915 / / / / / / / / / RIGHT BUNDLE BRANCH BLOCK 1155 0.976 0.925 0.925 0.967 0.911 0.976 0.932 0.922 0.931 0.946 0.842 0.912 RIGHT VENTRICULAR HYPERTROPHY 136 0.913 0.833 0.859 / / / / / / / / / SEPTAL INFARCT 1423 0.947 0.894 0.880 / / / / / / / / / SINUS BRADYCARDIA 119 0.950 0.972 0.955 0.967 0.938 0.989 0.996 0.966 0.946 0.983 0.935 0.963 SINUS RHYTHM 16785 0.923 0.936 0.751 / / / 0.977 0.942 0.920 0.963 0.926 0.913 SINUS TACHYCARDIA 826 0.994 0.976 0.987 0.988 0.946 0.996 0.986 0.957 0.968 0.979 0.932 0.944 SUPRAVENTRICULAR TACHYCARDIA 27 0.995 0.976 0.959 / / / / / / / / / VENTRICULAR TACHYCARDIA 24 0.987 0.905 0.993 / / / / / / / / / WITH 1ST DEGREE AV BLOCK 82 0.911 0.661 0.921 0.962 0.925 0.992 0.943 0.783 0.911 0.930 0.735 0.893 WITH QRS WIDENING 45 0.618 0.414 0.742 / / / / / / / / / WOLFF-PARKINSON-WHITE 18 0.919 0.747 0.982 / / / / / / / / /
Table 2: Performance of other ECG deep learning model and ECGFounder on chronic kidney disease ECG test set. For ECGFounder, this is internal validation; for the Ouyang’s model, this is external validation.
Models ECGFounder Ouyang[13]
Class Count 12--lead AUC 1--lead AUC 12--lead AUC 1--lead AUC
Chronic kidney disease 13,990 0.698 0.672 0.611 0.622
Chronic kidney disease, stage 2 (mild) 514 0.590 0.590 0.528 0.549
Chronic kidney disease, stage 3 (moderate) 5708 0.664 0.643 0.563 0.588
Chronic kidney disease, stage 4 (severe) 834 0.714 0.691 0.611 0.632
Chronic kidney disease, stage 5 533 0.734 0.708 0.584 0.611
End stage renal disease 6401 0.795 0.725 0.642 0.634
4 Discussion
In this study, we have developed and demonstrated the generalizability and robust diagnostic capabilities of ECGFounder, a universal foundation model for electrocardiogram analysis. Trained on the largest ECG dataset to date, HEEDB, which encompasses over ten million ECGs from more than one million unique patients across 150 primary diagnostic categories including normal ECGs, arrhythmias, conduction blocks, myocardial infarctions, and cardiac hypertrophy. We developed and validated a deep learning model that consistently outperforms other ECG models. Our findings on both internal and external test sets highlight the substantial clinical diagnostic value and generalizability of our model. Furthermore, we enhanced the model’s performance on single-lead ECGs through a novel data augmentation method based on the cardiac axis. The internal validation results for arrhythmia diagnosis using single-lead ECGs demonstrated exceptional performance, broadening the prospects for the model’s application in mobile health. Moreover, by leveraging a fine-tuned pre-trained model, ECGFounder effectively adapts to a wide range of downstream tasks, significantly enhancing the diagnosis of other diseases such as chronic kidney disease and coronary heart disease.
ECGFounder enhances diagnostic performance by learning to identify ECG features associated with cardiovascular diseases, which are typically diagnosed based on specific waveform patterns and rhythm characteristics like the elevated ST segments of myocardial infarction and the irregular fluctuations of atrial fibrillation. These features involve anomalies in cardiac electrical activity, appearing significantly different from normal ECG waveforms. Upon training, ECGFounder can recognize these disease-related waveform patterns and rhythms, accurately diagnosing corresponding cardiovascular conditions. As observed in Tables LABEL:tab:internal and Figure 2, ECGFounder matches or even exceeds the performance of cardiology experts on the internal review test sets. Furthermore, ECGFounder’s sensitivity exceeds that of the average cardiology expert, indicating its ability to more accurately capture subtle signs of cardiovascular diseases that may be overlooked by human experts.
Previous research has demonstrated that deep learning models can support clinical ECG analysis and achieve good performance.[11] However, existing models lack a universal clinical diagnostic capability for ECGs. Firstly, the training datasets used by existing models are not large or diverse enough, which can lead to overfitting and poor performance on new data, thus limiting their generalizability.[33] Additionally, the lack of demographic diversity in the training datasets means these models may perform poorly for certain demographic groups. This could lead to biases when the models are applied to data from different demographic backgrounds, affecting the fairness and accuracy of the models. Secondly, the labels for most model training datasets are derived from manual annotations by cardiology experts, which is time-consuming and labor-intensive. This limits the number of ECGs available for training. Also, because cardiology experts typically use a unified annotation system, the richness of the dataset labels is not very high, often only including common ECG abnormalities and omitting many important but rare diagnostic labels.[35] Thirdly, due to the training methods used, these models do not support both 12-lead and single-lead ECGs. To address these issues, we propose a foundation model for ECGs that supports both 12-lead and single-lead ECGs. In Tables 3 and 4, we observed that ECGFounder ranks first in average performance across various external tests. The other baseline models, including S12L-ECG, CTN, and ECG-SE-ResNet, have previously achieved the best performance in AI-ECG models.[24, 43, 31] S12L-ECG was trained using supervised learning on the CODE-ECG dataset, which includes 2 million ECGs with 6 common types of arrhythmia diagnoses.[31] CTN and ECG-SE-ResNet were trained on the PhysioNet Challenge-2020 dataset, which includes 60,000 ECGs covering 27 common ECG diagnoses.[43, 24] In this paper, we demonstrate that by training on a larger and more diverse ECG dataset, a scalable foundation model can be developed to further improve the diagnosis of cardiovascular diseases and surpass previous baseline models.
Despite the effectiveness of ECGFounder in detecting various cardiovascular diseases, there are still challenges to be addressed. Firstly, as most data used to develop ECGFounder come from U.S. cohorts, this limits the diversity and representativeness of the data. Different regions of the world may have unique ECG characteristics, such as race and regional-specific rhythm variations, requiring the model to handle data from diverse backgrounds and populations. Secondly, although AI models can provide high-accuracy diagnostic results, their decision-making process is often a ’black box,’ which might encounter trust and acceptance issues in clinical applications. Therefore, developing explainable AI models, allowing doctors to understand the model’s decision-making process, is key to advancing the use of ECG AI models. Lastly, some clinically relevant information, such as patients’ lifestyles and medical histories, which could serve as effective covariates in cardiovascular disease research, have not yet been included in the model. We suggest that future work should involve incorporating a larger volume of ECG data from different regions and ethnicities, adding demographic information of patients as model inputs, and developing more explainable AI models, enabling doctors to better understand the decision processes and outcomes of the models. At the same time, some of the latest natural language processing methods have shown great potential in handling ECG annotations from experts, allowing for a more effective use of the clinical knowledge embedded in these annotations for learning.[20]
5 Conclusion
In this work, we proposed and validated the efficacy of ECGFounder in adapting to a wide range of cardiovascular diagnostic applications, demonstrating its high performance and versatility across various downstream tasks as a foundational ECG model. By overcoming the limitations related to ECG data and labeling diversity, as well as training methods, the ECG foundation model confirms its potential to transform the standard of care in cardiology and to provide real-time, accurate cardiac assessments in diverse clinical settings.
Supplement
Appendix A Details of Dataset
In this section, the comparison and split methods of the dataset are introduced in detail.
Table LABEL:tb:data is a supplement to Figure 1.c in the main text. The table compares the size and label types of ECG datasets used in multiple existing works. It can be observed that HEEDB has far more data sizes and label types than other datasets.
Table 3: Comparison with ECG dataset of HEEDB and existing works.
# ECG # Patients # Classes # Lead
Stanford-ECG [11] 91,232 53,549 12 1
PTB-XL [38] 21,837 18,885 71 12
PhysioNet Challenge 2020 [26] 66,361 Unknown 27 12
CODE [31] 2,322,513 1,676,384 6 12
Tongji-ECG [42] 180,112 70,692 21 12
UCSF-ECG [16] 992,748 365,009 38 12
HEEDB (Ours) [19] 10,771,552 1,818,247 150 12
Table 4: Age, working year and education level of the cardiologists participating in the comparison with ECGFounder.
Member Working year Age Education level
Student Cardiologist 1 2 25 M.D. Candidate
Student Cardiologist 2 3 27 M.D. Candidate
Student Cardiologist 3 1 24 M.D. Candidate
Expert Cardiologist 1 25 50 Master
Expert Cardiologist 2 27 48 Master
Table 5: Details on fine-tuning dataset split.
Dataset Task type Train Valid Test
HEEDB classification 6,683,587 835,448 834,926
MIMIC-ECG-NTproBNP classification 47,084 5,884 5,893
MIMIC-ECG-CKD classification 213,926 26,749 26,741
MIMIC-ECG-LVEF classification 60,333 7,539 7,541
MIMIC-ECG-CHD classification 213,926 26,749 26,741
MIMIC-ECG-Age classification 640,027 80,015 80,004
MIMIC-ECG-Sex classification 640,027 80,015 80,004
MIMIC-ECG-NTproBNP regression 47,084 5,884 5,893
MIMIC-ECG-LVEF regression 60,333 7,539 7,541
MIMIC-ECG-Age regression 640,027 80,015 80,004
Deepbeat(PPG) classification 466,904 51,878 17,617
Appendix B Details of Labels
Table 6: Filtered 150 meaningful label list of HEEDB development set.
Label Count
ABNORMAL ECG 5916870
NORMAL SINUS RHYTHM 5440634
NORMAL ECG 1912774
SINUS RHYTHM 1583298
SINUS BRADYCARDIA 1188526
ATRIAL FIBRILLATION 1108003
SINUS TACHYCARDIA 964612
OTHERWISE NORMAL ECG 927256
LEFT AXIS DEVIATION 880984
PREMATURE VENTRICULAR COMPLEXES 834752
BORDERLINE ECG 697210
RIGHT BUNDLE BRANCH BLOCK 684092
SEPTAL INFARCT 662652
LEFT ATRIAL ENLARGEMENT 637167
NONSPECIFIC T WAVE ABNORMALITY 578612
LOW VOLTAGE QRS 528770
PREMATURE ATRIAL COMPLEXES 498470
ANTERIOR INFARCT 380691
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 377899
PREMATURE SUPRAVENTRICULAR COMPLEXES 333798
LEFT BUNDLE BRANCH BLOCK 318791
NONSPECIFIC T WAVE ABNORMALITY NOW EVIDENT IN 299977
NONSPECIFIC T WAVE ABNORMALITY NO LONGER EVIDENT IN 275433
T WAVE INVERSION NOW EVIDENT IN 264182
LATERAL INFARCT 263295
NONSPECIFIC ST ABNORMALITY 260714
LEFT VENTRICULAR HYPERTROPHY 257036
T WAVE INVERSION NO LONGER EVIDENT IN 251761
WITH RAPID VENTRICULAR RESPONSE 243049
QT HAS SHORTENED 221554
QT HAS LENGTHENED 216377
FUSION COMPLEXES 198828
ATRIAL FLUTTER 198007
MARKED SINUS BRADYCARDIA 183847
WITH SINUS ARRHYTHMIA 182984
NONSPECIFIC ST AND T WAVE ABNORMALITY 177718
LEFT ANTERIOR FASCICULAR BLOCK 154028
RIGHT AXIS DEVIATION 153561
ECTOPIC ATRIAL RHYTHM 151104
UNDETERMINED RHYTHM 148020
ANTEROSEPTAL INFARCT 136656
RIGHTWARD AXIS 130331
ST NOW DEPRESSED IN 128118
WITH SHORT PR 126460
WITH MARKED SINUS ARRHYTHMIA 124492
ST NO LONGER DEPRESSED IN 113896
INVERTED T WAVES HAVE REPLACED NONSPECIFIC T WAVE ABNORMALITY IN 109365
NON-SPECIFIC CHANGE IN ST SEGMENT IN 109261
NONSPECIFIC T WAVE ABNORMALITY HAS REPLACED INVERTED T WAVES IN 109217
JUNCTIONAL RHYTHM 108731
ELECTRONIC ATRIAL PACEMAKER 107890
ABERRANT CONDUCTION 103597
ELECTRONIC VENTRICULAR PACEMAKER 96125
T WAVE INVERSION LESS EVIDENT IN 94837
ANTEROLATERAL INFARCT 92527
WITH REPOLARIZATION ABNORMALITY 91043
RSR’ OR QR PATTERN IN V1 SUGGESTS RIGHT VENTRICULAR CONDUCTION DELAY 90752
T WAVE INVERSION MORE EVIDENT IN 90476
WIDE QRS RHYTHM 89366
WITH PREMATURE VENTRICULAR OR ABERRANTLY CONDUCTED COMPLEXES 88985
RIGHT ATRIAL ENLARGEMENT 64829
INFERIOR INFARCT 64044
INCOMPLETE LEFT BUNDLE BRANCH BLOCK 63540
VOLTAGE CRITERIA FOR LEFT VENTRICULAR HYPERTROPHY 62822
OR DIGITALIS EFFECT 62191
BIFASCICULAR BLOCK 59790
ST NO LONGER ELEVATED IN 58002
WITH SLOW VENTRICULAR RESPONSE 56777
ST ELEVATION NOW PRESENT IN 56383
PREMATURE ECTOPIC COMPLEXES 55818
LEFT POSTERIOR FASCICULAR BLOCK 55139
T WAVE AMPLITUDE HAS DECREASED IN 44131
WITH A COMPETING JUNCTIONAL PACEMAKER 42770
RIGHT SUPERIOR AXIS DEVIATION 40107
BIATRIAL ENLARGEMENT 40051
VENTRICULAR-PACED RHYTHM 39485
ATRIAL-PACED RHYTHM 39290
T WAVE AMPLITUDE HAS INCREASED IN 38556
WITH QRS WIDENING 37944
WITH 1ST DEGREE AV BLOCK 36157
PROLONGED QT 34564
WITH PROLONGED AV CONDUCTION 34345
RIGHT VENTRICULAR HYPERTROPHY 34095
WITH QRS WIDENING AND REPOLARIZATION ABNORMALITY 33211
ATRIAL-SENSED VENTRICULAR-PACED RHYTHM 32599
AV SEQUENTIAL OR DUAL CHAMBER ELECTRONIC PACEMAKER 31180
PULMONARY DISEASE PATTERN 31111
ACUTE MI / STEMI 30735
INFERIOR-POSTERIOR INFARCT 29775
NONSPECIFIC INTRAVENTRICULAR CONDUCTION DELAY 28423
PREMATURE VENTRICULAR AND FUSION COMPLEXES 26524
IN A PATTERN OF BIGEMINY 25765
AV DUAL-PACED RHYTHM 25056
SUPRAVENTRICULAR TACHYCARDIA 18945
VENTRICULAR-PACED COMPLEXES 17459
WIDE QRS TACHYCARDIA 17029
RSR’ PATTERN IN V1 16953
ST LESS DEPRESSED IN 15676
VENTRICULAR TACHYCARDIA 15603
EARLY REPOLARIZATION 15177
ST MORE DEPRESSED IN 14426
ANTEROLATERAL LEADS 14209
ELECTRONIC DEMAND PACING 14089
RBBB AND LEFT ANTERIOR FASCICULAR BLOCK 11593
LATERAL INJURY PATTERN 10815
BIVENTRICULAR PACEMAKER DETECTED 10686
SUSPECT UNSPECIFIED PACEMAKER FAILURE 10498
WOLFF-PARKINSON-WHITE 10264
WITH VENTRICULAR ESCAPE COMPLEXES 10240
INFERIOR INJURY PATTERN 10033
CONSIDER RIGHT VENTRICULAR INVOLVEMENT IN ACUTE INFERIOR INFARCT 9978
ST ELEVATION HAS REPLACED ST DEPRESSION IN 9585
NONSPECIFIC INTRAVENTRICULAR BLOCK 9258
MASKED BY FASCICULAR BLOCK 9074
PEDIATRIC ECG ANALYSIS 8707
BLOCKED 8529
WITH UNDETERMINED RHYTHM IRREGULARITY 8365
LEFTWARD AXIS 7690
WITH 2ND DEGREE SA BLOCK MOBITZ I 7380
ACUTE 6857
ABNORMAL LEFT AXIS DEVIATION 6708
WITH COMPLETE HEART BLOCK 6505
NO P-WAVES FOUND 6384
ST LESS ELEVATED IN 5406
WITH RETROGRADE CONDUCTION 5186
ST MORE ELEVATED IN 4982
JUNCTIONAL BRADYCARDIA 4651
WITH VARIABLE AV BLOCK 4556
ANTERIOR INJURY PATTERN 4412
WITH JUNCTIONAL ESCAPE COMPLEXES 4077
ACUTE MI 3867
ACUTE PERICARDITIS 3745
POSTERIOR INFARCT 3657
IDIOVENTRICULAR RHYTHM 3609
WITH 2ND DEGREE SA BLOCK MOBITZ II 2912
R IN AVL 2723
SINUS/ATRIAL CAPTURE 2603
AV DUAL-PACED COMPLEXES 2516
INFEROLATERAL INJURY PATTERN 2204
RBBB AND LEFT POSTERIOR FASCICULAR BLOCK 1922
ANTEROLATERAL INJURY PATTERN 1804
ATRIAL-PACED COMPLEXES 1655
WITH SINUS PAUSE 1518
BIVENTRICULAR HYPERTROPHY 1453
ABNORMAL RIGHT AXIS DEVIATION 1314
SUPRAVENTRICULAR COMPLEXES 1291
WITH 2ND DEGREE AV BLOCK MOBITZ I 1152
WITH 2:1 AV CONDUCTION 1080
WITH AV DISSOCIATION 1030
MULTIFOCAL ATRIAL TACHYCARDIA 1016
Appendix C Details of Model
Figure 4: The detailed architecture plot of the ECGFounder, indicating the dimensions of key components of the model.
Table 7: Hyperparameters for pre-training.
Hyperparameters Values
Batch size 1024
Learning rate 1e-3
Minimal learning rate 1e-6
Learning rate scheduler Cosine
Optimizer AdamW
Weight decay 0.1
Total epochs 20
Table 8: Hyperparameters for fine-tuning.
Hyperparameters Values
Batch size 256
Learning rate 1e-4
Minimal learning rate 1e-6
Learning rate scheduler ReduceLROnPlateau
Optimizer AdamW
Weight decay 0.1
Total epochs 30
Details of Results
Table 9: Performance of deep learning model and cardiologists on the internal test set
Model Cardiologists
Class Count AUC Sens Spec Sens Spec
ANTERIOR INFARCT 2 0.908 1.000 0.906 0.750 0.999
ATRIAL FIBRILLATION 59 0.996 1.000 0.972 0.798 0.981
ATRIAL FLUTTER 19 0.993 1.000 0.986 0.568 0.990
ATRIAL-PACED RHYTHM 3 0.991 1.000 0.988 0.833 0.995
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 4 0.893 0.959 0.898 0.563 0.977
INFERIOR INFARCT 1 0.998 1.000 0.998 1.000 0.997
LATERAL INFARCT 2 0.998 1.000 0.998 0.250 0.999
LEFT BUNDLE BRANCH BLOCK 11 1.000 1.000 0.998 0.600 0.932
NORMAL SINUS RHYTHM 364 0.969 0.934 0.926 0.916 0.862
PREMATURE ATRIAL COMPLEXES 19 0.997 1.000 0.963 0.579 0.989
PREMATURE VENTRICULAR COMPLEXES 40 0.981 0.975 0.889 0.850 0.996
RIGHT AXIS DEVIATION 12 0.996 1.000 0.992 0.594 0.979
RIGHT BUNDLE BRANCH BLOCK 50 0.984 0.940 0.969 0.775 0.970
SINUS BRADYCARDIA 41 0.995 1.000 0.989 0.717 0.994
SINUS RHYTHM 191 0.970 0.970 0.926 0.916 0.862
SINUS TACHYCARDIA 58 0.996 1.000 0.945 0.821 0.990
VENTRICULAR TACHYCARDIA 9 0.903 1.000 0.882 0.875 0.996
VENTRICULAR-PACED RHYTHM 20 0.988 0.950 0.940 0.725 0.997
WITH 1ST DEGREE AV BLOCK 15 0.864 0.667 0.831 0.633 0.962
WITH SINUS ARRHYTHMIA 6 0.976 1.000 0.945 0.542 0.979
Figure 5: ROC curves on the internal test set of the deep learning model, compared with cardiologists
Table 10: Performance of other ECG deep learning model and ECGFounder on external test set (single-lead ECG) PhysioNet Challenge-2017
Models ECGFounder Stanford[11]
Class AUC Sens Spec F1 F1
SINUS RHYTHM 0.975 0.961 0.983 0.941 0.910
ATRIAL FIBRILLATION 0.957 0.924 0.979 0.756 0.840
Figure 6: AUROC of single-lead ECG downstream tasks.
Figure 7: Regression scatter of ECG downstream tasks.
Appendix D Data link
Appendix E Model weights link
Will be released on https://bdsp.io/
Appendix F Code link