### v3

Abstract
BACKGROUND. Artificial intelligence (AI) has demonstrated significant potential in ECG analysis and cardiovascular disease assessment. Recently, foundation models have played a remarkable role in advancing medical AI, bringing benefits such as efficient disease diagnosis and cross-domain knowledge transfer. The development of an ECG foundation model holds the promise of elevating AI-ECG research to new heights. However, building such a model faces several challenges, including insufficient database sample sizes and inadequate generalization across multiple domains. Additionally, there is a notable performance gap between single-lead and multi-lead ECG analyses.
METHODS. We introduced an ECG Foundation Model (ECGFounder), a general-purpose model that leverages real-world ECG annotations from cardiology experts to broaden the diagnostic capabilities of ECG analysis. ECGFounder was trained on over 10 million ECGs with 150 label categories from the Harvard-Emory ECG Database, enabling comprehensive cardiovascular disease diagnosis through ECG analysis. The model is designed to be both an effective out-of-the-box solution, and a to be fine-tunable for downstream tasks, maximizing usability. Importantly, we extended its application to lower rank ECGs, and arbitrary single-lead ECGs in particular. ECGFounder is therefore applicable to supporting various downstream tasks in mobile and remote monitoring scenarios.
RESULTS. Experimental results demonstrate that ECGFounder achieves expert-level performance on internal validation sets, with AUROC exceeding 0.95 for eighty diagnoses. It also shows strong classification performance and generalization across various diagnoses on external validation sets. When fine-tuned, ECGFounder outperforms baseline models in demographic analysis, clinical event detection, and cross-modality cardiac rhythm diagnosis, surpassing baseline methods by 3 to 5 points in AUROC.
CONCLUSIONS. The ECG foundation model offers an effective solution, allowing it to generalize across a wide range of tasks. By enhancing existing cardiovascular diagnostics and supporting integration into mobile monitoring systems, it contributes to the advancement of the cardiovascular AI community and enables management of cardiac conditions.
keywords:
[…]
1 Introduction
The Electrocardiogram (ECG) is a crucial diagnostic tool that measures and records the electrical activity of the heart using electrodes placed on the skin.1 ECG recordings are essential for diagnosis and monitoring of cardiac health conditions. However, fully interpreting an ECG is complex and requires significant training and time. A typical ECG expert undergoes nearly 10 years of training, including medical school, internal medicine residency, and specialized ECG training.2 In recent years, the advent of deep learning together with efforts to assemble relatively large databases of ECGs have seen some interesting progress in the field, extending ECG analysis beyond the traditional medical domains 3, 4, 5, 6. However, due to the lack of large-scale publicly available ECG databases with diverse diagnostic information, developing a general-purpose AI-ECG model remains a challenging task. Existing models are often confined to specific diagnostic tasks and datasets. This challenge highlights the practical limitations of training an ECG model from scratch on small-scale datasets, making it difficult to extend the model to real-world ECG analysis.
Foundation models, with their strong generalization capabilities, provide a promising approach for enhancing the performance of AI-ECG models through transfer learning. Recently, these models have achieved significant advancements in the field of medical AI. In retinal disease diagnosis, the RETFound model, through pre-training on a large number of retinal images, has achieved excellent performance across various clinical diagnostic tasks.7 In computational pathology, the UNI model was trained on a vast amount of whole-slide images, reaching expert-level performance in multiple cancer diagnostic tasks.8 In these studies, foundation models are defined as large-scale AI models trained on extensive datasets, capable of adapting to a wide range of downstream tasks. Specifically, they meet the following criteria: 1) the pre-training dataset is large in scale, 2) the model has an enormous number of parameters, and 3) it can perform a wide range of downstream tasks.9
There have been several claims in the literature of developing foundation models for the ECG.10, 11 However, due to limitations of existing ECG databases in terms of sample size, patient numbers, the variety of diagnoses, and importantly, the demographics of the patients, these models have yet to address the challenges of diversity across regions, ethnic groups, and diagnostic variations.12 Moreover, to qualify as a foundation model, the trained model must be capable of generalizing to multiple domains outside of the initial training paradigm. 13 Additionally, current ECG models have significant performance degradation in single-lead ECGs compared to multi-lead ECGs.14
In this work, we propose the ECG Foundation Model, ECGFounder, which is capable of diagnosing 150 types of cardiac abnormalities, leveraging over ten million clinically annotated real-world ECGs spanning all existing ECG classifications. This represents the largest and most comprehensive ECG foundation model to date. Moreover, we demonstrate how the model is applicable to a wide range of tasks in varying domains, a fundamental requirement for foundation models.We provide the medical machine learning community with an accessible ECG foundation model that offers an effective out-of-the-box solution and fine-tuning capabilities. Compared to traditional AI-ECG models trained from scratch, ECGFounder offers a novel approach that achieves superior performance through fine-tuning. This advancement has the potential to drive the future development of AI-ECG technology.
To address the inherent challenges of incomplete annotations in real-world data, we introduce a novel method for pre-processing and training on these annotations, ensuring robust performance even under sub-optimal conditions. Moreover, by training the single-lead ECG model based on lead augmentation, we are able to maintain high diagnostic performance on single-lead ECGs as well. We validate the model’s performance on both internal and external test sets, where it consistently matches expert-level diagnoses. In downstream task fine-tuning, we demonstrate ECGFounder’s versatility in addressing various tasks, including demographics detection, clinical event detection, and cross-modality diagnosis (Figure 1b). Specifically, we evaluate ECGFounder on 12 clinical tasks, such as ECG age regression and classification, sex detection, chronic kidney disease (CKD) detection, chronic heart disease (CHD) detection, regression and abnormal classification of left ventricular ejection fraction (LVEF), regression and abnormal classification of NT-proBNP, and atrial fibrillation detection based on photoplethysmography (PPG). The results of these downstream tasks highlight ECGFounder’s potential as a foundational model for the further development of AI-ECG models.
The trained model and labeled data will be publicly released upon publication through the Brain Data Science Platform (bdsp.io), huggingface and PhysioNet, promoting further research and development in this field.
Figure 1: ECGFounder is a general ECG encoder based on the RegNet architecture. (a) ECGFounder was trained on 10,771,552 ECGs with 150 types of ECG diagnostic labels. Due to the presence of missing diagnoses in real-world expert annotations, we implemented positive label augmentation by modifying the loss function of the pre-training method. For more details, see the Methods section. (b) ECGFounder was applied to 12 clinical downstream tasks, covering ECG diagnosis, demographics inference, clinical event detection, and cross-modality diagnosis. In comparison with baseline methods, ECGFounder achieved state-of-the-art performance across all tasks. (c) A comparison of the HEEDB dataset used by ECGFounder with other large ECG datasets. (d) Data used in model development and validation for the ECG diagnosis and downstream tasks. PPG., photoplethysmography; LVEF., left ventricular ejection fraction; CKD., chronic kidney disease; CHD., chronic heart disease.
2 Methods
2.1 Datasets and pre-processing
Our dataset, the Harvard-Emory ECG Database (HEEDB), is currently the largest open-access ECG dataset, containing 10,771,552 expert-annotated ECGs from 1,818,247 unique subjects.15 These ECGs are predominantly 10-second, 12-lead clinical ECGs. The dataset includes annotations from cardiologists and ECG technicians paired with the ECGs. These annotations consist of discrete text reports that primarily describe the morphology, rhythm, and diagnostic information of the ECGs. Experts used the Marquette 12SL ECG Analysis Program (GE Healthcare) version 4 to assist with annotations, which provides waveform parameters for doctors’ reference.16 The program offers many diagnostic categories, allowing doctors to simply click on corresponding category labels on the computer, avoiding manual entry of diagnostic statements. To extract descriptive information about ECGs from these discrete labels, we utilized regular expressions to parse the annotations, tallying each independent label. In total, there were 287 independent phrases. After reviewing with doctors, we removed phrases that were irrelevant to ECG descriptions while retaining meaningful phrases. Ultimately, we defined 150 meaningful labels, including rich diagnostic information as well as specific rhythm and morphological descriptions (See Supplement for more details about labels).
Our external validation data comprised three large ECG databases (DB): The CODE-test DB 17, the PTB-XL DB 18, and PhysioNet Challenge-2017 DB 19. The CODE-test DB is composed of ECG records from 827 patients across 811 municipalities in Brazil, collected by the Telehealth Network of Minas Gerais (TNMG). There are six common arrhythmia labels in this ECG DB, annotated by several experienced ECG experts.
The PTB-XL dataset contains 21,837 clinical ECGs from 18,885 patients in Germany. Each ECG is a 10-second, 12-lead recording. The labels were reviewed and verified by two physicians. This dataset is currently one of the best publicly accessible ECG collections, both in terms of the number of samples and the quality of the labels.18
The PhysioNet Challenge-2017 is a large single-lead ECG dataset. The ECG recordings were collected using the AliveCor device.19 The training set includes 8,528 single-lead ECG recordings, with durations ranging from 9 to 60 seconds, while the test set contains 3,658 ECG recordings of similar lengths. The dataset requires classification of ECG recordings into normal rhythm, atrial fibrillation rhythm, other rhythms, and noise.
Additionally, we utilized the MIMIC-IV-ECG DB to fine-tune our model for various downstream tasks. The MIMIC-IV-ECG dataset is part of the MIMIC series, focusing on the collection and analysis of ECG data.20, 21 MIMIC-IV-ECG originates from real clinical settings at Beth Israel Deaconess Medical Center (BIDMC) in Boston, USA, and contains 800,035 clinical ECGs from 161,352 patients treated in the intensive care unit (ICU). Moreover, ECG recordings in the dataset can be matched with the electronic health records of patients in the MIMIC-ED, allowing ECG data to be associated with specific conditions. Here, we linked several clinical downstream tasks, such as age, sex, chronic kidney disease (CKD), chronic heart disease (CHD), left ventricular ejection fraction (LVEF) and NT-proBNP with ECG data to explore the performance improvement of the fine-tuned ECGFounder model in detecting other diseases and clinical events. More details about the split and labels of MIMIC-IV-ECG DB can be found in Supplement.
To explore ECGFounder’s generalization capabilities on other similar physiological signals, we fine-tuned and evaluated it using the DeepBeat dataset, a PPG-based atrial fibrillation (AF) detection dataset.22 The dataset includes over 500,000 labeled signals from more than 100 individuals.
For ECG preprocessing, we applied linear interpolation to resample ECG frequencies to 500Hz. We used a high-pass filter with a cutoff frequency of 1 Hz to suppress residual baseline drift and a second-order 30 Hz Butterworth low-pass filter to reduce high-frequency noise. A 50/60 Hz notch filter was utilized to eliminate electrical interference. From ECG records longer than 10 seconds, we extracted 10-second windows in sequence. If a sequence was less than 10 seconds, we applied zero padding. We normalized all signals using the mean and standard deviation of each individual signal segment before inputting them into the model. If a signal was missing, we assigned it a value of 0.
2.2 Model architecture
To establish an ECG foundation model, we require an architecture that is tailored for ECG, capable of learning generalizable representations from large-scale ECG datasets. The growth in the size of ECG data and the increase in the number of leads mean that the model must not only consider temporal information of the data but also spatial relationships, i.e., interactions between different leads and the overall pattern of cardiac electrical activity. This is crucial for the ECG foundation model to be applicable in real clinical ECG scenarios, to mitigate the impact of non-uniform ECG durations and missing leads in the training dataset.
Considering these factors, we built our model architecture on top of the RegNet architecture.23, 24 This structure begins with a stage-wise network design where each stage consists of a set number of blocks and channels that scale with network depth. This is beneficial for us to expand and design the blocks and channels of the ECG foundation model. Unlike traditional uniform scaling across layers, RegNet employs a quantized linear model to predict optimal widths and depths, ensuring efficient performance across a range of model sizes. The initial layers focus on capturing low-level features with fewer channels, which gradually increase as the network deepens, thus optimizing computational efficiency and model capacity. Following this, the model utilizes a series of bottleneck blocks that combine group convolutions and channel-wise attention mechanisms, enhancing the richness of representation in both temporal and spatial dimensions while controlling model complexity. This tailored configuration makes the model suitable for ECG data. More details about the model architecture can be found in the Supplement.
2.3 Training with real-world annotations: noisy, imbalanced, positive unlabeled
Unlike conventional ECG diagnostic models that typically use single-label classification methods, we employed multi-label classification during the training phase of the ECG foundation model. This approach aligns more closely with clinical practice, where an ECG diagnosis often consists of multiple diagnostic labels. For example, an abnormal ECG diagnosis might be something like normal sinus rhythm | premature ventricular complexes | premature ectopic complexes. Additionally, multi-label diagnostics better meet the application needs of clinicians. During training, utilizing a multi-label classification approach provides our model with rich semantic value and facilitates generalization to different annotation systems. However, the nature of multi-label annotations means that it is almost impossible to achieve perfect multi-label data, especially in cases like ECG where there are many diagnostic categories. Generally speaking, if a cardiology expert can annotate three classes simultaneously, it is considered excellent. However, the actual number of ECG diagnostic labels far exceeds three classes.25 To address the challenge brought by incomplete ECG annotations, we introduce positive unlabeled (PU) learning method.
PU learning is defined as when positive samples are present in the data but unlabeled, but the labeled positive samples are correct, meaning that the unlabeled samples are not necessarily negative examples.26 Typically, PU learning leads to a severe imbalance in the predicted probability distribution: an ECG usually contains a few positive diagnostic labels, with the remaining labels being treated as negative, leading to a severe positive-negative imbalance where the negative samples far outnumber the positive samples for each label. The inherent scarcity of positive examples, combined with the missing labels, exacerbates the imbalance, making it more challenging to learn from the positive examples. When using conventional loss functions, such as the Binary Cross-Entropy (BCE) loss function, the model tends to learn from simpler samples, namely the true negative (TN) samples, while the more challenging samples, namely the false negative (FN) samples, are harder to fit. When the model is biased towards learning TN samples, its predicted probabilities tend to skew towards 0 rather than 1, which diminishes its ability to predict positive samples and weakens its potential to detect clinically important abnormalities.
To enhance the model’s ability to fit representations of positive samples, we improved the multi-label classification loss function, enabling the model to correct missing labels by adjusting the weights of positive and negative samples. Typically, for missing labels, a well-trained multi-label model’s predicted probability should be close to 1 rather than 0. Therefore, by applying a smaller weight to the loss of negative labels with predicted probabilities close to 1, we can mitigate the impact of missing labels. The loss function is given by:
$\mathcal{L}=-(\gamma-p)p^{2}.$ (1)
Here, $\gamma$ is a hyperparameter of the model and $p$ is the predicted probability of model. In this case, it is set to $\gamma=1.5$, which we find optimally balances the weights, allowing the model to learn good representations of both positive and negative samples.
Our model training used AdamW to minimize the loss function, with an initial learning rate set to 0.001. The learning rate decayed by a factor of 10 every 5 epochs. The trainable temperature parameter was initialized to 0. Training spanned a maximum of 20 epochs, with early stopping based on validation loss. We used a batch size of 1024.
2.4 Training a single-lead ECG model based on leads augmentation
The development of portable and wearable ECG devices in recent years has revolutionized continuous cardiac monitoring, offering a non-invasive method for real-time health assessment. Beyond the conventional diagnosis of cardiac arrhythmias, another critical challenge in this field is accurately detecting and interpreting ECG axis deviation on single-lead ECGs from wearable devices (typically lead I), which can significantly impact the diagnosis of various cardiac abnormalities. For instance, left axis deviation can provide additional insights into diagnoses like left ventricular hypertrophy, left bundle branch block, left anterior fascicular block, pre-excitation syndromes, and inferior myocardial infarction.27 To address this issue, we developed a novel training method utilizing lead-augmented wearable ECG models. By systematically enhancing standard 12-lead ECG data, we simulate various clinical scenarios of axis inversion, thereby enhancing the model’s robustness and versatility. Understanding the relationship between ECG vectors and leads is crucial for this method. The cardiac electrical activity generates a vector representation of cardiac signals, captured by different ECG leads placed on the body. Each lead provides a unique perspective of the cardiac electrical axis, offering a comprehensive view when combined. The standard 12-lead ECG system includes limb leads (I, II, III, aVR, aVL, aVF) and precordial leads (V1-V6), with each lead representing a specific projection of the cardiac electrical vector.28 Specifically, we primarily utilize the ECG signals from limb leads. Based on the angular position of each limb lead’s axis relative to the heart, we consider lead I as the center of the semicircle (i.e., 0°) and calculate the signals from all leads around the semicircle (i.e., from -90° to 90°), thereby obtaining six additional leads (aVL, -aVR, II, -III, aVF, -aVF). We then trained a model for wearable ECG devices, extracting the lead I ECG from the HEEDB 12-lead data and randomly incorporating one of the remaining six augmented leads into the model for training with a 50% probability. This ensures that the model can learn arrhythmia features from lead I data and axis deviation from the additional six leads. Additionally, we have scaled down the model’s parameter size to optimize for wearable devices with limited computational resources.
2.5 Fine-tuning on ECG foundation model
When adapting to specific ECG downstream tasks, we need to retain the parameters of the base model and discard the initial classification linear layer. The number of classes in the downstream task determines the number of neurons needed in the final layer of the new linear layer. The training objective is to generate classification outputs that match the labels.
The total training period is 30 epochs, with a learning rate adjustment strategy that utilizes the scheduler. After every epoch, the scheduler monitors the specified metric, and if the performance does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 0.1. The learning rate reduction is triggered based on the maximization of the monitored metric. This approach ensures a dynamic adjustment of the learning rate depending on the training progress. After training in each epoch, the model is evaluated on the validation set.8 The model weights with the highest AUC on the validation set are saved as a checkpoint for future evaluation.
2.6 Clinical validation
To validate and compare the performance of our model, we established a committee consisting of three experienced ECG physicians to annotate the internal test set. We developed an ECG annotation system for physicians to use, and set up 20 types of labels including cardiac rate abnormalities, conduction blocks, myocardial dilation, myocardial infarction, and ECG morphologies, with several sub-labels under each category. The Supplement displays the complete list of label types. After independent annotation by the committee, a consensus determination was made; ECGs that did not reach consensus were removed, providing an expert standard for model evaluation. ECGs that the committee could not interpret or agree upon were eliminated from our test dataset.
Additionally, to compare the diagnostic accuracy between the model and experts, five additional ECG physicians were involved in this study. They were given specific instructions on how to use the system. The Supplement shows the doctors’ ages, levels of experience, and education. Each doctor was required to independently annotate each ECG from the internal test set. The annotations from these physicians were then compared to the model’s results.
The evaluation of the model was conducted by calculating accuracy, the area under the Receiver Operating Characteristic curve (AUROC), sensitivity, and specificity, each with bilateral 95% confidence intervals (CI).
3 Results
Our dataset comprises 10,771,552 expert-annotated ECGs recorded from 1,818,247 unique subjects using 12-lead ECG machines, covering the period from the 1990s to present, with a recording duration of 10 seconds. The average age of patients is 60.6 years, with 48.7% being female.15 The development dataset includes 7,519,035 ECGs from 1,725,643 patients, and the test dataset includes 834,926 ECGs from 265,818 patients. The committee internal test set includes 523 ECGs recorded from 2023, corresponding to 523 unique patients. We obtained a total of 150 labels describing the ECGs, including heart rate abnormalities, conduction abnormalities, cardiac hypertrophy, myocardial infarction, myocardial ischemia, pacemaker devices, and ECG waveform morphology. Using these labels, we trained two deep learning ECG models separately: one for standard 12-lead ECGs and the other designed for single-lead wearable device ECGs (lead I). We evaluated the performance of both models on internal and external test sets.
We first conducted large-scale validation of ECGFounder using an internal test dataset comprising 834,926 ECGs, to assess its consistency with diagnoses confirmed by clinically recognized cardiologists (Figure S3). ECGFounder demonstrated an impressive AUC of at least 0.950 across 82 diagnostic categories and at least 0.900 across 112 categories. Overall, the AUCs were particularly high for categories such as rhythm disorders, conduction abnormalities, cardiac hypertrophy, and myocardial infarction. In contrast, the AUCs were lower for categories involving waveform changes, which may be attributed to the annotation habits of certain physicians who tend to place less emphasis on documenting waveform changes in their diagnoses.
We validated the deep learning model for 12-lead ECGs using the committee internal test set. The algorithm’s average AUC score for diagnosing all 20 classifications was 0.968 (95% CI, 0.965–0.969), sensitivity was 0.971 (95% CI, 0.970–0.973), and specificity was 0.937 (95% CI, 0.935–0.938). When comparing the model with cardiologists, our model achieved an overall average F1 score of 0.677, outperforming the cardiologists’ average F1 score of 0.640. The model’s performance was compared with cardiologists’ performance across 20 diagnostic categories (Table S7 in Supplement). When comparing the model’s ROC curve with the true positive rate and false positive rate of cardiologists, the model outperformed the average performance of cardiologists for most labels (Figure S2 in Supplement).
In the external test set experiments for 12-lead ECGs, we evaluated the performance of our model and other models on the CODE-test and PTB-XL datasets. On CODE-test, our model achieved an average AUC of 0.981, outperforming other baseline models such as CinC2020-CTN, which had an average AUC of 0.963, and 0.980 (S12L-ECG) (Table 2). On PTB-XL, because the other two models can only diagnose arrhythmias and cannot complete other diagnostic classifications in the broader classes on PhysioNet Challenge-2020, we validated this dataset only on the class which the models have. Our model achieved an average F1 score of 0.732 and an average AUC of 0.924 (Table 2). These results indicate that our model generalizes to different regions, hospitals, and patients.
In the internal test set experiments focused on single-lead ECGs, the model demonstrated excellent performance in rhythm-type diagnoses(Figure S3). Notably, it achieved an AUC above 0.95 for common heart rate abnormalities such as normal sinus rhythm, sinus bradycardia, sinus tachycardia, marked sinus bradycardia, sinus arrhythmia, marked sinus arrhythmia, and atrial fibrillation. These diagnoses can be reliably identified from single-lead ECGs. Additionally, the model achieved an AUC above 0.8 for diagnoses including premature ventricular complexes, premature supraventricular complexes, pacemaker, 1st degree AV block, branch block, fascicular block, and chamber enlargement. It is worth noting that for some of these diagnoses, this is the first time such exceptional performance has been observed using single-lead ECGs. In the diagnosis of myocardial infarctions, the model also showed good diagnostic performance for lateral infarct, anterolateral infarct, acute MI, and STEMI. In ECG research, these diagnoses are associated with shifts in the heart’s electrical axis. This suggests that data augmentation methods based on the electrical axis are effective for training single-lead ECG models.
In the external test set experiments on single-lead ECG devices, our model accurately classified sinus rhythm and atrial fibrillation. The model’s performance, as shown in the table, achieved an AUC of 0.975 and 0.957 for normal sinus rhythm and atrial fibrillation, respectively (Table S8 in Supplement). This indicates that the model achieves excellent performance on portable and wearable ECGs in real-world scenarios.
Table 1: Performance of ECGFounder and cardiologists on the committee test set
ECGFounder Cardiologists
Class Count AUC Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 2 0.908 1.000 0.906 0.408 0.750 0.999 0.625
ATRIAL FIBRILLATION 59 0.996 1.000 0.972 0.866 0.798 0.981 0.796
ATRIAL FLUTTER 19 0.993 1.000 0.986 0.759 0.568 0.990 0.514
ATRIAL-PACED RHYTHM 3 0.991 1.000 0.988 0.500 0.833 0.995 0.657
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 4 0.893 1.000 0.698 0.506 0.563 0.977 0.299
INFERIOR INFARCT 1 0.998 1.000 0.998 0.667 1.000 0.997 0.725
LATERAL INFARCT 2 0.998 1.000 0.998 0.667 0.250 0.999 0.250
LEFT BUNDLE BRANCH BLOCK 11 1.000 1.000 0.998 0.909 0.600 0.932 0.371
NORMAL SINUS RHYTHM 364 0.969 0.934 0.926 0.952 0.916 0.862 0.930
PREMATURE ATRIAL COMPLEXES 19 0.997 1.000 0.963 0.679 0.579 0.989 0.613
PREMATURE VENTRICULAR COMPLEXES 40 0.981 0.975 0.889 0.600 0.850 0.996 0.898
RIGHT AXIS DEVIATION 12 0.996 1.000 0.992 0.800 0.594 0.979 0.489
RIGHT BUNDLE BRANCH BLOCK 50 0.984 0.940 0.969 0.847 0.775 0.970 0.762
SINUS BRADYCARDIA 41 0.995 1.000 0.989 0.938 0.717 0.994 0.791
SINUS RHYTHM 191 0.970 0.970 0.926 0.971 0.916 0.862 0.930
SINUS TACHYCARDIA 58 0.996 1.000 0.945 0.683 0.821 0.990 0.833
VENTRICULAR TACHYCARDIA 9 0.903 1.000 0.882 0.635 0.875 0.996 0.678
VENTRICULAR-PACED RHYTHM 20 0.988 0.950 0.940 0.559 0.725 0.997 0.756
WITH 1ST DEGREE AV BLOCK 15 0.864 0.667 0.831 0.287 0.633 0.962 0.560
WITH SINUS ARRHYTHMIA 6 0.976 1.000 0.945 0.308 0.542 0.979 0.319
Figure 2: (a) Results of NTproBNP classification, NTproBNP regression, CKD classification, LVEF classification, LVEF regression and CHD classification tasks of ECGFounder and other baseline models for clinical events detection tasks. (b) Results of age classification over 65 years old, age regression and sex classification tasks of ECGFounder and other baseline models for demographics detection tasks. (c) Results of PPG atrial fibrillation detection tasks.
We next validated the performance of ECGFounder in transfer learning. Our model was fine-tuned and evaluated on six downstream tasks using supervised learning on the MIMIC-IV-ECG dataset, resulting in both a 12-lead ECG model and a single-lead ECG model. The fine-tuning results are shown in Figures 2. As shown, ECGFounder outperforms the baseline methods in every downstream task. Specifically, it achieves 2 to 3 percentage points higher performance than ECG-SimCLR and 4 to 6 points higher than ECG-ResNet in age, sex, NTproBNP, CKD, and CHD detection. Additionally, to compare with the previously published CKD study 29, we created an independent test set consisting of individual patients from the MIMIC-IV-ECG dataset for validation. As shown in the results (Table 3), our model surpasses the previously studied model in all six classifications for both 12-lead and single-lead ECGs.
Table 2: Performance of other ECG deep learning model and ECGFounder on external test set CODE-test and PTB-XL.17, 18 In particular, for the baseline method S12L-ECG, the CODE-test is the internal test set.
Dataset CODE-test Models ECGFounder S12L-ECG(Internal)17 CTN30 ECG-SE-ResNet31 Count AUC Sens Spec AUC Sens Spec AUC Sens Spec AUC Sens Spec SINUS BRADYCARDIA 16 0.967 1.000 0.955 0.955 0.938 0.996 0.965 0.987 0.942 0.932 0.995 0.937 ATRIAL FIBRILLATION 13 0.999 1.000 0.996 0.963 0.769 1.000 0.966 0.972 0.969 0.976 0.980 0.970 SINUS TACHYCARDIA 37 0.989 0.974 0.970 0.977 0.973 0.997 0.972 0.958 0.943 0.976 0.950 0.957 RIGHT BUNDLE BRANCH BLOCK 34 0.989 0.971 0.971 0.995 1.000 0.995 0.989 0.946 0.986 0.965 0.949 0.947 LEFT BUNDLE BRANCH BLOCK 30 0.998 1.000 0.996 1.000 1.000 1.000 0.961 0.988 0.942 0.971 0.949 0.976 WITH 1ST DEGREE AV BLOCK 28 0.949 0.864 0.957 0.989 0.929 0.995 0.930 0.862 0.925 0.958 0.856 0.922 Dataset PTB-XL Models ECGFounder S12L-ECG17 CTN30 ECG-SE-ResNet31 ANTERIOR INFARCT 36 0.635 0.816 0.447 / / / / / / / / / ANTEROLATERAL INFARCT 29 0.945 0.837 0.918 / / / / / / / / / ANTEROSEPTAL INFARCT 2415 0.945 0.891 0.882 / / / / / / / / / ATRIAL FIBRILLATION 1514 0.993 0.969 0.982 0.925 0.776 0.989 0.953 0.935 0.915 0.950 0.946 0.907 ATRIAL FLUTTER 73 0.993 0.973 0.969 / / / 0.948 0.962 0.937 0.941 0.966 0.956 ELECTRONIC ATRIAL PACEMAKER 298 0.909 0.935 0.745 / / / / / / / / / INFERIOR INFARCT 276 0.851 0.825 0.732 / / / / / / / / / LATERAL INFARCT 1066 0.916 0.819 0.859 / / / / / / / / / LEFT ANTERIOR FASCICULAR BLOCK 167 0.972 0.922 0.919 / / / 0.965 0.882 0.914 0.968 0.819 0.901 LEFT ATRIAL ENLARGEMENT 427 0.799 0.742 0.704 / / / / / / / / / LEFT BUNDLE BRANCH BLOCK 539 0.989 0.969 0.965 0.983 0.963 0.991 0.912 0.932 0.933 0.931 0.917 0.930 LEFT POSTERIOR FASCICULAR BLOCK 18 0.840 0.605 0.904 / / / / / / / / / LEFT VENTRICULAR HYPERTROPHY 2419 0.900 0.860 0.788 / / / / / / / / / LOW VOLTAGE QRS 182 0.581 0.690 0.423 / / / 0.605 0.605 0.814 0.593 0.655 0.741 N-SPECIFIC INTRAVENTRICULAR BLOCK 78 0.766 0.339 0.932 / / / 0.788 0.487 0.906 0.776 0.699 0.793 NORMAL ECG 987 0.887 0.952 0.639 / / / / / / / / / PREMATURE VENTRICULAR COMPLEXES 1146 0.987 0.965 0.959 / / / 0.963 0.955 0.921 0.937 0.939 0.880 QT HAS LENGTHENED 119 0.931 0.897 0.827 / / / 0.930 0.881 0.957 0.930 0.901 0.930 RIGHT ATRIAL ENLARGEMENT 99 0.959 0.869 0.915 / / / / / / / / / RIGHT BUNDLE BRANCH BLOCK 1155 0.976 0.925 0.925 0.967 0.911 0.976 0.932 0.922 0.931 0.946 0.842 0.912 RIGHT VENTRICULAR HYPERTROPHY 136 0.913 0.833 0.859 / / / / / / / / / SEPTAL INFARCT 1423 0.947 0.894 0.880 / / / / / / / / / SINUS BRADYCARDIA 119 0.950 0.972 0.955 0.967 0.938 0.989 0.996 0.966 0.946 0.983 0.935 0.963 SINUS RHYTHM 16785 0.923 0.936 0.751 / / / 0.977 0.942 0.920 0.963 0.926 0.913 SINUS TACHYCARDIA 826 0.994 0.976 0.987 0.988 0.946 0.996 0.986 0.957 0.968 0.979 0.932 0.944 SUPRAVENTRICULAR TACHYCARDIA 27 0.995 0.976 0.959 / / / / / / / / / VENTRICULAR TACHYCARDIA 24 0.987 0.905 0.993 / / / / / / / / / WITH 1ST DEGREE AV BLOCK 82 0.911 0.661 0.921 0.962 0.925 0.992 0.943 0.783 0.911 0.930 0.735 0.893 WITH QRS WIDENING 45 0.618 0.414 0.742 / / / / / / / / / WOLFF-PARKINSON-WHITE 18 0.919 0.747 0.982 / / / / / / / / /
Table 3: Performance of other ECG deep learning model and ECGFounder on chronic kidney disease ECG test set. For ECGFounder, this is internal validation; for the Ouyang’s model, this is external validation.
Models ECGFounder Ouyang29
Class Count 12--lead AUC 1--lead AUC 12--lead AUC 1--lead AUC
Chronic kidney disease 13,990 0.698 0.672 0.611 0.622
Chronic kidney disease, stage 2 (mild) 514 0.590 0.590 0.528 0.549
Chronic kidney disease, stage 3 (moderate) 5708 0.664 0.643 0.563 0.588
Chronic kidney disease, stage 4 (severe) 834 0.714 0.691 0.611 0.632
Chronic kidney disease, stage 5 533 0.734 0.708 0.584 0.611
End stage renal disease 6401 0.795 0.725 0.642 0.634
4 Discussion
In this study, we have developed and demonstrated the generalizability and robust diagnostic capabilities of ECGFounder, a universal foundation model for electrocardiogram analysis. Trained on the largest ECG dataset to date, HEEDB, which encompasses over ten million ECGs from more than one million unique patients across 150 primary diagnostic categories including normal ECGs, arrhythmias, conduction blocks, myocardial infarctions, and cardiac hypertrophy. We developed and validated a deep learning model that consistently outperforms other ECG models. Our findings on both internal and external test sets highlight the substantial clinical diagnostic value and generalizability of our model. Furthermore, we enhanced the model’s performance on single-lead ECGs through a novel data augmentation method based on the cardiac axis. The internal validation results for arrhythmia diagnosis using single-lead ECGs demonstrated exceptional performance, broadening the prospects for the model’s application in mobile health. Moreover, by leveraging a fine-tuned pre-trained model, ECGFounder effectively adapts to a wide range of downstream tasks, significantly enhancing the diagnosis of other diseases such as chronic kidney disease and coronary heart disease.
ECGFounder enhances diagnostic performance by learning to identify ECG features associated with cardiovascular diseases, which are typically diagnosed based on specific waveform patterns and rhythm characteristics like the elevated ST segments of myocardial infarction and the irregular fluctuations of atrial fibrillation. These features involve anomalies in cardiac electrical activity, appearing significantly different from normal ECG waveforms. Upon training, ECGFounder can recognize these disease-related waveform patterns and rhythms, accurately diagnosing corresponding cardiovascular conditions. As observed in Table S7 in Supplement and Figure S3, ECGFounder matches or even exceeds the performance of cardiology experts on the internal review test sets. Furthermore, ECGFounder’s sensitivity exceeds that of the average cardiology expert, indicating its ability to more accurately capture subtle signs of cardiovascular diseases that may be overlooked by human experts.
Previous research has demonstrated that deep learning models can support clinical ECG analysis and achieve good performance.32, 33, 6 However, existing models lack a universal clinical diagnostic capability for ECGs. Firstly, the training datasets used by existing models are not large or diverse enough, which can lead to overfitting and poor performance on new data, thus limiting their generalizability.34 Additionally, the lack of demographic diversity in the training datasets means these models may perform poorly for certain demographic groups. This could lead to biases when the models are applied to data from different demographic backgrounds, affecting the fairness and accuracy of the models. Secondly, the labels for most model training datasets are derived from manual annotations by cardiology experts, which is time-consuming and labor-intensive. This limits the number of ECGs available for training. Also, because cardiology experts typically use a unified annotation system, the richness of the dataset labels is not very high, often only including common ECG abnormalities and omitting many important but rare diagnostic labels.35 Thirdly, due to the training methods used, these models do not support both 12-lead and single-lead ECGs. To address these issues, we propose a foundation model for ECGs that supports both 12-lead and single-lead ECGs. In Tables 3 and 4, we observed that ECGFounder ranks first in average performance across various external tests. The other baseline models, including S12L-ECG, CTN, and ECG-SE-ResNet, have previously achieved the best performance in AI-ECG models.30, 31, 17 S12L-ECG was trained using supervised learning on the CODE-ECG dataset, which includes 2 million ECGs with 6 common types of arrhythmia diagnoses.17 CTN and ECG-SE-ResNet were trained on the PhysioNet Challenge-2020 dataset, which includes 60,000 ECGs covering 27 common ECG diagnoses.31, 30 In this paper, we demonstrate that by training on a larger and more diverse ECG dataset, a scalable foundation model can be developed to further improve the diagnosis of cardiovascular diseases and surpass previous baseline models.
Despite the effectiveness of ECGFounder in detecting various cardiovascular diseases, there are still challenges to be addressed. Firstly, as most data used to develop ECGFounder come from U.S. cohorts, this limits the diversity and representativeness of the data. Different regions of the world may have unique ECG characteristics, such as race and regional-specific rhythm variations, requiring the model to handle data from diverse backgrounds and populations. Secondly, although ECG foundation models can provide high-accuracy diagnostic results, their decision-making process is often a ’black box,’ which might encounter trust and acceptance issues in clinical applications. Therefore, developing explainable AI models, allowing doctors to understand the model’s decision-making process, is key to advancing the use of ECG AI models. Lastly, some clinically relevant information, such as patients’ lifestyles and medical histories, which could serve as effective covariates in cardiovascular disease research, have not yet been included in the model. We suggest that future work should involve incorporating a larger volume of ECG data from different regions and ethnicities, adding demographic information of patients as model inputs, and developing more explainable AI models, enabling doctors to better understand the decision processes and outcomes of the models. At the same time, some of the latest natural language processing methods have shown great potential in handling ECG annotations from experts, allowing for a more effective use of the clinical knowledge embedded in these annotations for learning.36
5 Conclusion
In this work, we proposed and validated the efficacy of ECGFounder in adapting to a wide range of cardiovascular diagnostic applications, demonstrating its high performance and versatility across various downstream tasks as a foundational ECG model. By overcoming the limitations related to ECG data and labeling quality and diversity, as well as training methods, our ECG foundation model confirms its potential to transform the standard of care in cardiology and to provide real-time, accurate cardiac assessments in diverse clinical settings. We have provided open access to the model, as well as the code and data used to train it under an open access license. In this way, we invite the research community to continue build upon our model, and help advance the state-of-the-art.
Acknowledgements
This document is the result of the research project funded by the National Science Foundation and Emory University via an unrestricted gift.
Retrospective analysis of data for this project was conducted with waiver of informed consent under approved IRB protocols (BIDMC: 2022P000417; MGH: 2013P001024).
Funding: Dr. Westover was supported by grants from the NIH (RF1AG064312, RF1NS120947, R01AG073410, R01HL161253, R01NS126282, R01AG073598, R01NS131347, R01NS130119), and NSF (2014431). Dr. Shenda Hong was supported by the National Natural Science Foundation of China (62102008), Clinical Medicine Plus X - Young Scholars Project of Peking University, the Fundamental Research Funds for the Central Universities (PKU2024LCXQ030). For this work, Dr. Clifford was partially supported by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under NIH award number R01EB030362.
Disclosures: Dr. Westover is a co-founder, scientific advisor, and consultant to Beacon Biosignals and has a personal equity interest in the company. Dr. Clifford holds significant stock in AliveCor.
The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health, or the authors’ current and past employers and funding bodies.
We thank Dr. Xinxin Di, Dr. Kun Lu, Dr. Zhengkai Xue, Dr. Wenbo Dai, Dr. Jing Zhao, Dr. Hongqian Zhou for annotating the internal test set.
Supplement
S1. Q&A of the ECGFounder
What is the ECG foundation model?
The ECG foundation model is a large-scale, pre-trained neural network designed for tasks involving ECG (electrocardiogram) data. This model aims to capture and learn generalizable features from ECG signals, which can be fine-tuned or adapted for a wide range of downstream tasks like disease classification, arrhythmia detection, or patient-specific anomaly identification.
In our study, we propose the ECGFounder, which is being developed using the Harvard-Emory ECG Dataset. The model leverages this diverse and comprehensive dataset to create a robust feature extractor that can later be fine-tuned or used in transfer learning scenarios for specific diagnostic or research objectives. The ultimate goal is to have a powerful, generalized model that can perform various ECG-related tasks with minimal additional training on specific datasets.
What is the use of ECG foundation model?
We provide the medical AI community with a versatile ECG foundation model that serves as a comprehensive feature extractor and a highly adaptable base for transfer learning tasks. Compared to traditional models trained from scratch, this foundation model offers a generalized representation of ECG signals, enabling enhanced performance across diverse downstream tasks through fine-tuning. By leveraging large-scale pre-training on the Harvard-Emory ECG Dataset, the model captures robust and clinically relevant features, thus reducing the need for extensive labeled datasets for specific applications. This advancement holds significant potential to drive future innovations in AI-based ECG analysis, disease detection, and personalized healthcare solutions. And we offer an effective out-of-the-box solution and fine-tuning capabilities for the community to use it more easily.
How is the performance of the ECG foundation model?
The ECG foundation model demonstrates strong performance through several tasks, showcasing its high accuracy and generalization capabilities due to pre-training on large dataset like the Harvard-Emory ECG Dataset. It achieves superior results across various downstream tasks, including arrhythmia classification, disease diagnosis, and demographics detection, with fine-tuning further enhancing performance compared to models trained from scratch. The model effectively captures important ECG features, enabling precise differentiation of cardiac conditions with minimal domain-specific training. Additionally, its pre-trained architecture allows for efficient transfer learning, requiring fewer labeled samples and training epochs, thereby reducing time and computational costs. Its adaptability extends beyond disease classification to tasks like demographics and laboratory measurement detection, highlighting its versatility. Validation results consistently demonstrate that the foundation model outperforms baseline ECG models, achieving higher AUC, making it a robust and efficient tool for advancing AI applications in cardiology and beyond. The table S1 and S2 show the performance of ECG foundation model.
Table S1: Performance of classification tasks of the ECG foundation model.
Dataset Task type ECGFounder ECG-SimCLR ECG-ResNet
MIMIC-ECG-NTproBNP Laboratory measurement 0.8504 0.8133 0.8007
MIMIC-ECG-CKD Disease diagnosis 0.7459 0.7189 0.7093
MIMIC-ECG-LVEF Medical image metric 0.8674 0.8552 0.8543
MIMIC-ECG-CHD Disease diagnosis 0.7951 0.7631 0.7550
MIMIC-ECG-Age Demographics 0.8613 0.8394 0.8033
MIMIC-ECG-Sex Demographics 0.8983 0.8736 0.8412
Deepbeat(PPG) PPG AF detection 0.8203 0.7984 0.7950
Table S2: Performance of regression tasks of the ECG foundation model.
Dataset Task type ECGFounder ECG-SimCLR ECG-ResNet
MIMIC-ECG-NTproBNP Laboratory measurement 3.8468 4.8873 4.9893
MIMIC-ECG-LVEF Medical image metric 7.0117 7.0579 7.9514
MIMIC-ECG-Age Demographics 8.6501 9.0153 9.4004
How can researchers use the ECG Foundation Model?
The trained model and labeled data will be publicly released upon publication through the Brain Data Science Platform (bdsp.io), huggingface and PhysioNet.
S2. Details of Datasets
S2.1. Details of pre-training dataset
Table LABEL:tb:data is a supplement to Figure 1.c in the main text. The table compares the size and label types of ECG datasets used in multiple existing works. It can be observed that Harvard-Emory ECG Database(HEEDB) has far more data sizes and label types than other datasets. And table 2 describes the label types and corresponding quantities of all HEEDB datasets.
Table S3: Comparison with ECG dataset of HEEDB and existing works.
ECGs Patients Classes Lead
Stanford-ECG 32 91,232 53,549 12 1
PTB-XL 18 21,837 18,885 71 12
PhysioNet Challenge 2020 37 66,361 Unknown 27 12
CODE 17 2,322,513 1,676,384 6 12
Tongji-ECG 6 180,112 70,692 21 12
UCSF-ECG 33 992,748 365,009 38 12
HEEDB (Ours) 15 10,771,552 1,818,247 150 12
Table S4: Filtered 150 meaningful label list of HEEDB development set.
Label Count
ABNORMAL ECG 5916870
NORMAL SINUS RHYTHM 5440634
NORMAL ECG 1912774
SINUS RHYTHM 1583298
SINUS BRADYCARDIA 1188526
ATRIAL FIBRILLATION 1108003
SINUS TACHYCARDIA 964612
OTHERWISE NORMAL ECG 927256
LEFT AXIS DEVIATION 880984
PREMATURE VENTRICULAR COMPLEXES 834752
BORDERLINE ECG 697210
RIGHT BUNDLE BRANCH BLOCK 684092
SEPTAL INFARCT 662652
LEFT ATRIAL ENLARGEMENT 637167
NONSPECIFIC T WAVE ABNORMALITY 578612
LOW VOLTAGE QRS 528770
PREMATURE ATRIAL COMPLEXES 498470
ANTERIOR INFARCT 380691
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 377899
PREMATURE SUPRAVENTRICULAR COMPLEXES 333798
LEFT BUNDLE BRANCH BLOCK 318791
NONSPECIFIC T WAVE ABNORMALITY NOW EVIDENT IN 299977
NONSPECIFIC T WAVE ABNORMALITY NO LONGER EVIDENT IN 275433
T WAVE INVERSION NOW EVIDENT IN 264182
LATERAL INFARCT 263295
NONSPECIFIC ST ABNORMALITY 260714
LEFT VENTRICULAR HYPERTROPHY 257036
T WAVE INVERSION NO LONGER EVIDENT IN 251761
WITH RAPID VENTRICULAR RESPONSE 243049
QT HAS SHORTENED 221554
QT HAS LENGTHENED 216377
FUSION COMPLEXES 198828
ATRIAL FLUTTER 198007
MARKED SINUS BRADYCARDIA 183847
WITH SINUS ARRHYTHMIA 182984
NONSPECIFIC ST AND T WAVE ABNORMALITY 177718
LEFT ANTERIOR FASCICULAR BLOCK 154028
RIGHT AXIS DEVIATION 153561
ECTOPIC ATRIAL RHYTHM 151104
UNDETERMINED RHYTHM 148020
ANTEROSEPTAL INFARCT 136656
RIGHTWARD AXIS 130331
ST NOW DEPRESSED IN 128118
WITH SHORT PR 126460
WITH MARKED SINUS ARRHYTHMIA 124492
ST NO LONGER DEPRESSED IN 113896
INVERTED T WAVES HAVE REPLACED NONSPECIFIC T WAVE ABNORMALITY IN 109365
NON-SPECIFIC CHANGE IN ST SEGMENT IN 109261
NONSPECIFIC T WAVE ABNORMALITY HAS REPLACED INVERTED T WAVES IN 109217
JUNCTIONAL RHYTHM 108731
ELECTRONIC ATRIAL PACEMAKER 107890
ABERRANT CONDUCTION 103597
ELECTRONIC VENTRICULAR PACEMAKER 96125
T WAVE INVERSION LESS EVIDENT IN 94837
ANTEROLATERAL INFARCT 92527
WITH REPOLARIZATION ABNORMALITY 91043
RSR’ OR QR PATTERN IN V1 SUGGESTS RIGHT VENTRICULAR CONDUCTION DELAY 90752
T WAVE INVERSION MORE EVIDENT IN 90476
WIDE QRS RHYTHM 89366
WITH PREMATURE VENTRICULAR OR ABERRANTLY CONDUCTED COMPLEXES 88985
RIGHT ATRIAL ENLARGEMENT 64829
INFERIOR INFARCT 64044
INCOMPLETE LEFT BUNDLE BRANCH BLOCK 63540
VOLTAGE CRITERIA FOR LEFT VENTRICULAR HYPERTROPHY 62822
OR DIGITALIS EFFECT 62191
BIFASCICULAR BLOCK 59790
ST NO LONGER ELEVATED IN 58002
WITH SLOW VENTRICULAR RESPONSE 56777
ST ELEVATION NOW PRESENT IN 56383
PREMATURE ECTOPIC COMPLEXES 55818
LEFT POSTERIOR FASCICULAR BLOCK 55139
T WAVE AMPLITUDE HAS DECREASED IN 44131
WITH A COMPETING JUNCTIONAL PACEMAKER 42770
RIGHT SUPERIOR AXIS DEVIATION 40107
BIATRIAL ENLARGEMENT 40051
VENTRICULAR-PACED RHYTHM 39485
ATRIAL-PACED RHYTHM 39290
T WAVE AMPLITUDE HAS INCREASED IN 38556
WITH QRS WIDENING 37944
WITH 1ST DEGREE AV BLOCK 36157
PROLONGED QT 34564
WITH PROLONGED AV CONDUCTION 34345
RIGHT VENTRICULAR HYPERTROPHY 34095
WITH QRS WIDENING AND REPOLARIZATION ABNORMALITY 33211
ATRIAL-SENSED VENTRICULAR-PACED RHYTHM 32599
AV SEQUENTIAL OR DUAL CHAMBER ELECTRONIC PACEMAKER 31180
PULMONARY DISEASE PATTERN 31111
ACUTE MI / STEMI 30735
INFERIOR-POSTERIOR INFARCT 29775
NONSPECIFIC INTRAVENTRICULAR CONDUCTION DELAY 28423
PREMATURE VENTRICULAR AND FUSION COMPLEXES 26524
IN A PATTERN OF BIGEMINY 25765
AV DUAL-PACED RHYTHM 25056
SUPRAVENTRICULAR TACHYCARDIA 18945
VENTRICULAR-PACED COMPLEXES 17459
WIDE QRS TACHYCARDIA 17029
RSR’ PATTERN IN V1 16953
ST LESS DEPRESSED IN 15676
VENTRICULAR TACHYCARDIA 15603
EARLY REPOLARIZATION 15177
ST MORE DEPRESSED IN 14426
ANTEROLATERAL LEADS 14209
ELECTRONIC DEMAND PACING 14089
RBBB AND LEFT ANTERIOR FASCICULAR BLOCK 11593
LATERAL INJURY PATTERN 10815
BIVENTRICULAR PACEMAKER DETECTED 10686
SUSPECT UNSPECIFIED PACEMAKER FAILURE 10498
WOLFF-PARKINSON-WHITE 10264
WITH VENTRICULAR ESCAPE COMPLEXES 10240
INFERIOR INJURY PATTERN 10033
CONSIDER RIGHT VENTRICULAR INVOLVEMENT IN ACUTE INFERIOR INFARCT 9978
ST ELEVATION HAS REPLACED ST DEPRESSION IN 9585
NONSPECIFIC INTRAVENTRICULAR BLOCK 9258
MASKED BY FASCICULAR BLOCK 9074
PEDIATRIC ECG ANALYSIS 8707
BLOCKED 8529
WITH UNDETERMINED RHYTHM IRREGULARITY 8365
LEFTWARD AXIS 7690
WITH 2ND DEGREE SA BLOCK MOBITZ I 7380
ACUTE 6857
ABNORMAL LEFT AXIS DEVIATION 6708
WITH COMPLETE HEART BLOCK 6505
NO P-WAVES FOUND 6384
ST LESS ELEVATED IN 5406
WITH RETROGRADE CONDUCTION 5186
ST MORE ELEVATED IN 4982
JUNCTIONAL BRADYCARDIA 4651
WITH VARIABLE AV BLOCK 4556
ANTERIOR INJURY PATTERN 4412
WITH JUNCTIONAL ESCAPE COMPLEXES 4077
ACUTE MI 3867
ACUTE PERICARDITIS 3745
POSTERIOR INFARCT 3657
IDIOVENTRICULAR RHYTHM 3609
WITH 2ND DEGREE SA BLOCK MOBITZ II 2912
R IN AVL 2723
SINUS/ATRIAL CAPTURE 2603
AV DUAL-PACED COMPLEXES 2516
INFEROLATERAL INJURY PATTERN 2204
RBBB AND LEFT POSTERIOR FASCICULAR BLOCK 1922
ANTEROLATERAL INJURY PATTERN 1804
ATRIAL-PACED COMPLEXES 1655
WITH SINUS PAUSE 1518
BIVENTRICULAR HYPERTROPHY 1453
ABNORMAL RIGHT AXIS DEVIATION 1314
SUPRAVENTRICULAR COMPLEXES 1291
WITH 2ND DEGREE AV BLOCK MOBITZ I 1152
WITH 2:1 AV CONDUCTION 1080
WITH AV DISSOCIATION 1030
MULTIFOCAL ATRIAL TACHYCARDIA 1016
S2.2. Details of fine-tuning dataset
To evaluate the capability of transfer learning of ECGFounder, we developed 10 kinds of downstream datasets for fine-tuning.
For the MIMIC-ECG-NTproBNP dataset, we extracted data from the lab events section of MIMIC-IV-Hosp, obtaining NT-proBNP values matched with corresponding patients and admission IDs. According to clinically recognized laboratory standards, NT-proBNP values were categorized into two groups: normal and abnormal. This classification, along with the raw NT-proBNP values, was then matched to the corresponding ECGs of the patients for use in classification and regression tasks, respectively. Additionally, considering that abnormal NT-proBNP values can sometimes be excessively high, leading to non-convergence in the regression loss, we applied z-score normalization to these values. All normalized values were scaled to a range of 0 to 50.
For the MIMIC-ECG-CKD dataset, we extracted data from the diagnoses-ICD section of MIMIC-IV-Hosp, obtaining ICD codes matched with corresponding patients and admission IDs. Considering that the MIMIC database contains both ICD-9 and ICD-10 formats, we selected ICD-10 as the reference standard and converted all ICD codes to ICD-10. The ICD-10 codes for chronic kidney disease (CKD) include N182 (Chronic kidney disease, stage 2 [mild]), N183 (Chronic kidney disease, stage 3 [moderate]), N184 (Chronic kidney disease, stage 4 [severe]), N185 (Chronic kidney disease, stage 5), and N186 (End-stage renal disease). These codes were consolidated and treated collectively as indicators of CKD. If a patient has a positive diagnosis under any of these codes, they are considered to have CKD. This CKD classification information was then matched to the corresponding ECGs of the patients for use in the CKD classification task.
For the MIMIC-ECG-CHD dataset, we extracted ICD codes following the same procedure as described above. The ICD-10 codes for chronic heart disease (CHD) include all codes from I20 to I25. These codes were consolidated and treated collectively as indicators of CHD. If a patient has a positive diagnosis under any of these codes, they are considered to have CHD. This CHD classification information was then matched to the corresponding ECGs of the patients for use in the CHD classification task.
For the MIMIC-ECG-LVEF dataset, we extracted data from the discharge section of MIMIC-IV-Notes, obtaining echocardiogram notes and corresponding LVEF values matched with patients and admission IDs. According to clinically recognized echocardiogram diagnostic standards, an LVEF of 50% or higher is considered normal, while an LVEF below 50% is considered abnormal. This LVEF classification information, along with the raw LVEF values, was matched to the corresponding ECGs of the patients for use in both classification and regression tasks.
For the MIMIC-ECG-Age dataset, we extracted data from the patients section of MIMIC-IV-Hosp, obtaining age information matched with the corresponding patients. Ages were classified using a threshold of 65 years: individuals aged 65 or older were considered elderly, while those under 65 were not. This age classification, along with the raw age values, was matched to the corresponding ECGs of the patients for use in both classification and regression tasks.
For the MIMIC-ECG-Sex dataset, we extracted data from the patients section of MIMIC-IV-Hosp, obtaining sex information matched with the corresponding patients. This sex classification information was then matched to the corresponding ECGs of the patients for use in the sex classification task.
For the PPG Dataset, we utilized the DeepBeat-PPG dataset, which includes PPG data from 132 patients with atrial fibrillation (AF). This dataset provides publicly available validation and test sets. Here, we used the validation set as both the training and validation set, and the test set as a holdout test set. This approach ensures that patients do not overlap between the training and test sets, allowing for a fair performance comparison.
Table S5: Details on pre-training and fine-tuning dataset split.
Dataset Task type Train Valid Test
HEEDB(Pre-training) classification 6,683,587 835,448 834,926
MIMIC-ECG-NTproBNP classification 47,084 5,884 5,893
MIMIC-ECG-CKD classification 213,926 26,749 26,741
MIMIC-ECG-LVEF classification 60,333 7,539 7,541
MIMIC-ECG-CHD classification 213,926 26,749 26,741
MIMIC-ECG-Age classification 640,027 80,015 80,004
MIMIC-ECG-Sex classification 640,027 80,015 80,004
MIMIC-ECG-NTproBNP regression 47,084 5,884 5,893
MIMIC-ECG-LVEF regression 60,333 7,539 7,541
MIMIC-ECG-Age regression 640,027 80,015 80,004
Deepbeat(PPG) classification 466,904 51,878 17,617
S2.3. Details of cardiologists
Table S6 describes the personal information of 4 cardiologists participating in Cardiologist versus ECGFounder, including working year, gender, age and education level.
Table S6: Working year, gender, age and education level of the cardiologists participating in the comparison with ECGFounder.
Member Working year Gender Age Education level
Student Cardiologist 1 2 Male 25 M.D. Candidate
Student Cardiologist 2 3 Female 27 M.D. Candidate
Expert Cardiologist 1 27 Female 50 Master
Expert Cardiologist 2 25 Male 48 Master
S3. Details of Method
S3.1. Details of model architecture
Figure S1: The detailed architecture plot of the ECGFounder, indicating the dimensions of key components of the model.
In the input layer of ECGFounder, the model receives a multi-channel ECG signal represented by a size of 12×5000, indicating 12 channels and 5000 time steps. This input passes through a 1D convolutional layer with an input size of 12×5000, where 64 filters of size 64×64 are applied to produce a feature map with dimensions of 64×160. As the network progresses beyond the first layer, multiple convolutional blocks are introduced, with each block composed of convolution and pooling operations that gradually increase the network depth. Initially, these transformations reduce the feature map size from 64×160 to 32×80, eventually leading to an output size of 8×400, achieved through two repetitions of the convolutional block.
Subsequently, from the third to the fifth layer, the network continues to apply convolution and pooling, which results in further dimensional reduction while progressively extracting more intricate features. The third layer’s output remains at 400×400, with three repetitions of the convolution operation, while the fourth layer expands to produce a feature map of 1024×1024, repeated four times. Following these layers of convolution and feature extraction, the network transitions to a fully connected layer. Here, the extracted features from the preceding layers serve as input, and the fully connected operation generates either a single value or a categorical prediction based on the task at hand.
Notably, skip connections are included between layers to mitigate the vanishing gradient problem typically encountered in deep networks. These connections not only improve the training process but also enhance overall model performance by facilitating efficient information flow across the network.
S3.2. Details of training
The training process for ECGFounder involved distinct sets of hyperparameters for pre-training and fine-tuning stages, optimized to achieve robust model performance. During the pre-training phase, a large batch size of 1024 was employed alongside a relatively high learning rate of 1e-3, which was gradually reduced using a ReduceLROnPlateau scheduler to a minimum of 1e-5. The AdamW optimizer was utilized with a weight decay of 0.1 to enhance generalization, and training was conducted over a total of 20 epochs.
In the fine-tuning phase, the batch size was reduced to 256, and the learning rate was lowered to 1e-4 with a minimum threshold of 1e-6. The ReduceLROnPlateau scheduler was retained to adaptively manage the learning rate, while the AdamW optimizer and a weight decay of 0.1 continued to be employed. The total number of epochs was increased to 30, allowing the model to refine its performance on downstream tasks. This systematic approach to hyperparameter selection ensured the model’s adaptability and stability across various training stages.
S3.3. Details of ECG preprocessing
Here, we provide a more comprehensive description of our preprocessing pipeline . The description is as follows:
First, we used linear interpolation to resample each ECG signal to a fixed sampling rate of 500Hz. This step ensures that all ECG signals, regardless of their original sampling rate, have a consistent temporal resolution, facilitating downstream processing and analysis.
Next, we applied a high-pass filter with a cutoff frequency of 1Hz to suppress residual baseline drift. Baseline drift often arises from low-frequency variations such as respiration and electrode-skin impedance changes. By attenuating frequency components below 1Hz, this high-pass filter helps maintain the main morphological features of the ECG waveform without low-frequency noise.
After that, we employed a 2nd-order Butterworth low-pass filter with a cutoff frequency of 30Hz to reduce high-frequency noise. This filter preserves the primary frequency components of the ECG signal—typically within the 0–30Hz range—while attenuating higher frequencies that may be introduced by muscle artifacts or electrical interference.
Furthermore, to remove power-line interference, we used a notch filter at 50/60Hz. This notch filter specifically targets and attenuates the narrow frequency band at which power-line noise occurs, improving signal fidelity without excessively affecting the rest of the ECG spectrum.
Then, from any ECG recording longer than 10 seconds, we extracted consecutive 10-second segments. If a recording was shorter than 10 seconds, we applied zero-padding to the signal to reach the desired 10-second duration. This uniform window length ensures that each segment fed into the model has the same temporal dimension.
Before inputting these processed segments into the model, we performed Z-score normalization using each segment’s mean and standard deviation. Concretely, this involves subtracting the segment mean from each data point and dividing by the segment’s standard deviation, resulting in each segment having a mean of 0 and a standard deviation of 1. This step prevents signals with inherently larger amplitudes from overshadowing those with smaller amplitudes, making model training more stable.
Lastly, for any missing ECG signals, we simply assigned them a value of zero. In other words, these missing channels are represented as zero-filled data, ensuring a consistent input dimension for the model without introducing spurious numerical artifacts.
By following these steps, we effectively mitigate baseline drift, high-frequency noise, and power-line interference, while establishing a uniform sampling rate and window size for all signals, ultimately providing clean and standardized ECG segments for subsequent analysis and model training.
S4. Details of Experiments
S4.1. Details of ablation experiments
We provide a series of ablation experimental results to demonstrate the advantages and innovations of our method.
Here, we use the average AUC of the internal validation set and the average AUC of the CODE-test (external validation set) as our metrics, respectively.
The table are as follow:
Table S7: Effects of loss function.
Loss Internal External
BCE Loss 86.18 (86.12-86.24) 95.32 (95.31-95.33)
Focal Loss 88.67 (88.65-88.69) 96.73 (96.72-96.74)
PU Loss (Ours) 92.04 (92.00-92.07) 98.20 (98.15-98.25)
Table S8: Effects of hyperparameters of loss function.
Num of $gamma$ Internal External
0.5 91.18 (91.14-91.22) 97.22 (97.21-97.23)
1 91.77 (91.75-88.79) 97.37 (97.36-97.38)
1.5 (Ours) 92.04 (92.00-92.07) 98.20 (98.15-98.25)
2 91.96 (91.94-91.98) 98.12 (98.11-98.13)
Table S9: Effects of model architecture.
Model architecture Internal External
CNN-1d 87.45 (87.43-97.47) 93.21 (93.12-93.30)
ResNet-1d 91.85 (91.84-91.86) 97.64 (97.62-97.67)
Transformer-1d 91.53 (91.52-91.55) 97.78 (97.75-97.81)
RegNet-1d (Ours) 92.04 (92.01-92.07) 98.20 (98.18-98.23)
Table S10: Effects of 1-lead data argumentation.
Methods Internal(1-lead) External(1-lead)
Only 1-lead data 83.76 (83.62-83.90) 94.96 (94.83-95.09)
1-lead data argumentation (Ours) 86.92 (86.83-87.01) 96.37 (96.33-96.41)
Table S11: Effects of number of model parameters.
Model architecture Internal External
11.7M 89.62 (89.50-89.74) 96.67 (96.61-96.73)
25.6M 90.23 (90.04-90.41) 97.41 (97.30-97.52)
76.3M (Ours) 92.04 (92.01-92.07) 98.20 (98.18-98.23)
110M 91.87 (91.85-91.88) 98.34 (98.01-98.67)
Table S12: Effects of number of training set.
Model architecture Internal External
10K 86.78 (86.72-86.84) 91.11 (91.04-91.18)
100K 87.93 (87.93-87.93) 92.08 (92.03-92.13)
1M 89.95 (89.73-90.17) 96.21 (96.11-96.31)
3M 91.85 (91.80-91.89) 97.82 (97.72-97.91)
5M 91.93 (91.62-92.25) 97.78 (97.75-97.81)
10M (Ours) 92.04 (92.01-92.07) 98.20 (98.18-98.23)
S4.2. Details of linear probing
In order to explore the advantages of ECGFounder under few-shot fine-tuning and the generalization of features, we conducted linear probing experiments on an ECG self supervised learning benchmark.
For linear probing, we kept the backbone network frozen and only trained the randomly initialized parameters of the linear classifier. To explore the performance of our method under low-resource conditions, we conducted linear probing using 1%, 10%, and 100% of the training data for each task. We set the learning rate to $5\times 10^{-3}$ and trained for 100 epochs. All test results were obtained from the best validation model, rather than testing the model on the test set after each epoch and reporting the highest result. For all downstream tasks, we used the AUC as the evaluation metric. We set the random seed to 00 to ensure the reproducibility of all results.
Table S13: Linear probing results of ECGFounder and other SSL models. The best results are bolded, with gray indicating the second highest.
Method PTBXL-Super PTBXL-Sub PTBXL-Form PTBXL-Rhythm CPSC2018 CSN 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% SimCLR 38 63.41 69.77 73.53 60.84 68.27 73.39 54.98 56.97 62.52 51.41 69.44 77.73 59.78 68.52 76.54 59.02 67.26 73.20 TS-TCC 39 70.73 75.88 78.91 53.54 66.98 77.87 48.04 61.79 71.18 43.34 69.48 78.23 57.07 73.62 78.72 55.26 68.48 76.79 CLOCS 40 68.94 73.36 76.31 57.94 72.55 76.24 51.97 57.96 72.65 47.19 71.88 76.31 59.59 77.78 77.49 54.38 71.93 76.13 ST-MEM 41 61.12 66.87 71.36 54.12 57.86 63.59 55.71 59.99 66.07 51.12 65.44 74.85 56.69 63.32 70.39 59.77 66.87 71.36 HeartLang 42 78.94 85.59 87.52 64.68 79.34 88.91 58.70 63.99 80.23 62.08 76.22 90.34 60.44 66.26 77.87 57.94 68.93 82.49 ECG-FM 11 69.41 82.10 85.79 64.12 79.28 86.51 59.67 66.98 76.02 59.83 74.08 86.05 59.94 69.19 78.81 55.24 66.83 79.05 ECGFounder (Ours) 79.65 87.34 91.20 75.72 81.73 88.03 61.69 68.76 82.52 70.05 88.10 93.61 64.21 79.65 83.18 64.68 75.89 84.35
S4.3. Details of subgroup analysis
In order to explore the fairness of ECGFounder under various demographic conditions, we conducted subgroup analysis on ECGFounder. The table is as follows:
Table S14: Performance in age, sex, and race stratified populations.
Cohort Internal External
Less than 40 years old 91.85 (91.75-91.94) 98.21 (98.18-98.24)
40-60 years old 92.44 (92.39-92.48) 97.94 (97.92-97.96)
60-75 years old 91.58 (91.55-91.62) 97.78 (97.75-97.81)
Greater than 75 years old 92.03 (91.95-92.13) 98.21 (98.20-98.23)
Male 91.98 (91.92-92.14) 98.22 (98.18-98.26)
Female 92.10 (92.06-92.15) 98.16 (98.15-98.17)
White 92.72 (92.66-92.78) /
Asian 91.44 (91.40-92.48) /
African American 91.90 (91.86-91.94) /
American Indian 90.55 (90.46-90.64) /
Other races 92.11 (92.07-92.15) /
S5. Details of Results
S5.1. Details of internal committee validation
To further compare the performance of ECGFounder in clinical diagnosis with that of experts, we compared the performance of the model with that of 5 cardiologists. The results are shown in the figure S2.
Label Sensitivity Sensitivity_CI Specificity Specificity_CI
LATERAL INFARCT 1 (1.0, 1.0) 0.94 (0.92, 0.96)
INFERIOR INFARCT 1 (1.0, 1.0) 1 (0.99, 0.99)
PREMATURE ATRIAL COMPLEXES 0.95 (0.87, 1.0) 0.96 (0.95, 0.97)
SINUS TACHYCARDIA 0.93 (0.85, 1.0) 0.97 (0.96, 0.99)
RIGHT AXIS DEVIATION 1 (1.0, 1.0) 0.97 (0.96, 0.98)
ATRIAL FIBRILLATION 1 (1.0, 1.0) 0.97 (0.95, 0.98)
SINUS BRADYCARDIA 1 (1.0, 1.0) 0.98 (0.97, 0.99)
ATRIAL FLUTTER 0.91 (0.90, 0.94) 0.99 (0.98, 0.99)
ATRIAL-PACED RHYTHM 0.63 (0.32, 0.94) 1 (0.99, 1.0)
RIGHT BUNDLE BRANCH BLOCK 0.92 (0.88, 0.97) 0.97 (0.95, 0.98)
PREMATURE VENTRICULAR COMPLEXES 0.88 (0.79, 0.97) 0.95 (0.93, 0.97)
WITH SINUS ARRHYTHMIA 0.75 (0.67, 0.87) 0.96 (0.94, 0.96)
SINUS RHYTHM 0.98 (0.96, 0.99) 0.54 (0.48, 0.63)
NORMAL SINUS RHYTHM 0.94 (0.91, 0.96) 0.9 (0.89, 0.95)
ANTERIOR INFARCT 0 (0.0, 0.0) 0.93 (0.91, 0.95)
VENTRICULAR TACHYCARDIA 0 (0.0, 0.0) 1 (0.99, 1.0)
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.93 (0.85, 1.0) 0.99 (0.98, 1.0)
WITH 1ST DEGREE AV BLOCK 0.66 (0.62, 0.68) 1 (0.99, 1.0)
LEFT BUNDLE BRANCH BLOCK 1 (1.0, 1.0) 0.98 (0.97, 0.99)
VENTRICULAR-PACED RHYTHM 0.94 (0.91, 0.96) 1 (0.99, 1.0)
Label PPV PPV_CI NPV NPV_CI
LATERAL INFARCT 0.03 (0.0, 0.10) 1 (1.0, 1.0)
INFERIOR INFARCT 0.33 (0.0, 0.66) 1 (1.0, 1.0)
PREMATURE ATRIAL COMPLEXES 0.5 (0.39, 0.62) 1 (1.0, 1.0)
SINUS TACHYCARDIA 0.68 (0.51, 0.81) 1 (0.99, 1.0)
RIGHT AXIS DEVIATION 0.35 (0.27, 0.49) 1 (1.0, 1.0)
ATRIAL FIBRILLATION 0.75 (0.67, 0.87) 1 (1.0, 1.0)
SINUS BRADYCARDIA 0.81 (0.67, 0.89) 1 (1.0, 1.0)
ATRIAL FLUTTER 0.63 (0.32, 0.94) 1 (1.0, 1.0)
ATRIAL-PACED RHYTHM 0.68 (0.51, 0.81) 1 (0.99, 1.0)
RIGHT BUNDLE BRANCH BLOCK 0.78 (0.70, 0.88) 0.99 (0.98, 1.0)
PREMATURE VENTRICULAR COMPLEXES 0.63 (0.54, 0.74) 0.99 (0.98, 1.0)
WITH SINUS ARRHYTHMIA 0.15 (0.05, 0.25) 1 (0.99, 1.0)
SINUS RHYTHM 0.85 (0.81, 0.88) 0.91 (0.86, 0.98)
NORMAL SINUS RHYTHM 0.96 (0.95, 0.97) 0.84 (0.80, 0.92)
ANTERIOR INFARCT 0 (0.0, 0.0) 1 (1.0, 1.0)
VENTRICULAR TACHYCARDIA 0 (0.0, 0.0) 1 (0.99, 1.0)
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.13 (0.0, 0.34) 0.99 (0.99, 1.0)
WITH 1ST DEGREE AV BLOCK 0.5 (0.0, 1.0) 0.97 (0.96, 0.99)
LEFT BUNDLE BRANCH BLOCK 0.33 (0.10, 0.51) 1 (1.0, 1.0)
VENTRICULAR-PACED RHYTHM 0.71 (0.71, 1.0) 0.97 (0.96, 0.98)
Label AUROC AUROC_CI AUPRC AUPRC_CI
LATERAL INFARCT 1 (1.0, 1.0) 0.62 (0.57, 0.67)
INFERIOR INFARCT 1 (1.0, 1.0) 0.5 (0.39, 0.61)
PREMATURE ATRIAL COMPLEXES 1 (0.99, 1.0) 0.94 (0.86, 0.98)
SINUS TACHYCARDIA 1 (0.99, 1.0) 0.96 (0.90, 0.99)
RIGHT AXIS DEVIATION 1 (0.99, 1.0) 0.71 (0.29, 0.95)
ATRIAL FIBRILLATION 1 (0.99, 1.0) 0.95 (0.87, 0.98)
SINUS BRADYCARDIA 1 (0.99, 1.0) 0.9 (0.83, 0.99)
ATRIAL FLUTTER 0.99 (0.99, 1.0) 0.62 (0.34, 0.94)
ATRIAL-PACED RHYTHM 0.99 (0.99, 1.0) 0.35 (0.27, 0.49)
RIGHT BUNDLE BRANCH BLOCK 0.98 (0.97, 0.99) 0.81 (0.65, 0.89)
PREMATURE VENTRICULAR COMPLEXES 0.98 (0.97, 0.99) 0.81 (0.71, 0.91)
WITH SINUS ARRHYTHMIA 0.98 (0.96, 1.0) 0.37 (0.11, 0.49)
SINUS RHYTHM 0.97 (0.96, 0.99) 0.98 (0.95, 1.0)
NORMAL SINUS RHYTHM 0.97 (0.94, 0.98) 0.98 (0.97, 0.99)
ANTERIOR INFARCT 0.91 (0.86, 0.98) 0.66 (0.62, 0.68)
VENTRICULAR TACHYCARDIA 0.9 (0.89, 0.95) 0.54 (0.48, 0.63)
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.88 (0.79, 0.97) 0.33 (0.10, 0.51)
WITH 1ST DEGREE AV BLOCK 0.86 (0.80, 0.92) 0.22 (0.13, 0.41)
LEFT BUNDLE BRANCH BLOCK 1 (1.0, 1.0) 1 (1.0, 1.0)
VENTRICULAR-PACED RHYTHM 0.99 (0.98, 1.0) 0.74 (0.58, 0.89)
Table S15: The evaluation results of each cardiologist.
Cardiologist 1 Cardiologist 2
Class Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 1.000 1.000 1.000 1.000 0.996 0.500
ATRIAL FIBRILLATION 0.810 0.969 0.756 0.929 0.991 0.918
ATRIAL FLUTTER 0.818 0.982 0.521 0.909 0.992 0.544
ATRIAL-PACED RHYTHM 1.000 0.996 0.750 1.000 0.990 0.545
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.750 0.938 0.248 0.250 0.998 0.333
INFERIOR INFARCT 1.000 0.996 0.500 1.000 0.994 0.400
LATERAL INFARCT 0.000 1.000 0.000 0.000 0.998 0.000
LEFT BUNDLE BRANCH BLOCK 0.454 0.949 0.125 0.727 1.000 0.189
NORMAL SINUS RHYTHM 0.929 0.912 0.947 0.959 0.897 0.960
PREMATURE ATRIAL COMPLEXES 0.368 0.988 0.438 0.842 0.992 0.821
PREMATURE VENTRICULAR COMPLEXES 0.830 0.993 0.868 0.930 1.000 0.961
RIGHT AXIS DEVIATION 0.380 0.996 0.482 0.250 1.000 0.498
RIGHT BUNDLE BRANCH BLOCK 0.500 0.942 0.495 0.920 0.989 0.911
SINUS BRADYCARDIA 0.737 0.998 0.736 0.816 0.996 0.873
SINUS RHYTHM 0.929 0.912 0.947 0.959 0.897 0.960
SINUS TACHYCARDIA 0.929 0.998 0.945 0.714 0.998 0.816
VENTRICULAR TACHYCARDIA 1.000 0.998 0.667 0.500 1.000 0.566
VENTRICULAR-PACED RHYTHM 0.950 0.994 0.705 1.000 0.994 0.730
WITH 1ST DEGREE AV BLOCK 0.600 0.994 0.667 0.867 0.994 0.639
WITH SINUS ARRHYTHMIA 0.000 0.998 0.167 0.833 0.990 0.517
Table S16: The evaluation results of each cardiologist.
Cardiologist 3 Cardiologist 4
Class Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 1.000 1.000 1.000 0.000 1.000 0.000
ATRIAL FIBRILLATION 0.738 0.991 0.805 0.714 0.974 0.714
ATRIAL FLUTTER 0.091 0.998 0.409 0.455 0.992 0.493
ATRIAL-PACED RHYTHM 0.667 1.000 0.657 0.667 0.998 0.667
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 1.000 0.996 0.286 0.250 0.996 0.286
INFERIOR INFARCT 1.000 1.000 1.000 1.000 1.000 1.000
LATERAL INFARCT 0.000 1.000 0.000 1.000 1.000 1.000
LEFT BUNDLE BRANCH BLOCK 0.727 0.998 0.800 0.454 0.994 0.400
NORMAL SINUS RHYTHM 0.970 0.934 0.967 0.808 0.963 0.887
PREMATURE ATRIAL COMPLEXES 0.474 0.990 0.684 0.632 0.988 0.649
PREMATURE VENTRICULAR COMPLEXES 0.830 0.991 0.872 0.830 1.000 0.904
RIGHT AXIS DEVIATION 0.880 1.000 0.669 0.880 0.996 0.534
RIGHT BUNDLE BRANCH BLOCK 0.880 0.996 0.805 0.800 1.000 0.889
SINUS BRADYCARDIA 0.447 0.998 0.632 0.868 0.985 0.886
SINUS RHYTHM 0.970 0.934 0.967 0.808 0.963 0.887
SINUS TACHYCARDIA 0.750 0.994 0.852 0.893 0.970 0.746
VENTRICULAR TACHYCARDIA 1.000 1.000 0.800 1.000 0.998 0.667
VENTRICULAR-PACED RHYTHM 0.200 0.998 0.885 0.750 1.000 0.657
WITH 1ST DEGREE AV BLOCK 0.467 0.994 0.560 0.600 0.984 0.563
WITH SINUS ARRHYTHMIA 0.500 0.990 0.311 0.833 0.941 0.250
Figure S2: ROC curves on the internal test set of the deep learning model, compared with cardiologists
newpage
S5.2. Details of internal validation
Figure S3: Results of AUROC on 12-lead and single-lead ECG in internal test set.
S5.3. Details of external validation
We evaluated single-lead atrial fibrillation detection on an external validation set. As shown in the tableS17, the performance of our model exceeds the previous SOTA method.
Table S17: Performance of other ECG deep learning model and ECGFounder on external test set (single-lead ECG) PhysioNet Challenge-2017
Models ECGFounder Stanford32
Class AUC Sens Spec F1 F1
SINUS RHYTHM 0.975 0.961 0.983 0.941 0.910
ATRIAL FIBRILLATION 0.957 0.924 0.979 0.856 0.840
S5.4. Details of fine-tuning
Moreover, as shown in the figure S4 and S5, we evaluated several downstream tasks on single-lead ECGFounder thorough fine-tuning.
Figure S4: AUROC of single-lead ECG downstream tasks.
Figure S5: Regression scatter of single-lead ECG downstream tasks.
Also, as shown in the fig S6, we compare the performance of linear probing versus full fine-tuning, particularly focusing on the downstream classification tasks of CHD and LVEF. The detailed results clearly demonstrate that full fine-tuning consistently achieves better performance compared to linear probing, particularly for the ECGFounder model. Furthermore, the ECGFounder model shows superior performance compared to SIMCLR and random initialization under both fine-tuning methods, underscoring the strong representational capacity of ECGFounder.
Figure S6: Comparison of different fine-tuning methods, taking the classification of CHD and LVEF as examples.
Table S18: The PPV and prevalence of the fine-tuning task.
Task PPV(%) Prevalence(%)
Age 77.1 53.9
CHD 65 41.4
CKD 24.2 13.2
Lab 94.1 82.4
LVEF 94.1 76.9
Sex 83.7 51.7
Figure S7: 18 images arranged in a 3x6 grid.
S6. Data link
S7. Model weights link
S8. Code link